### v1

# Sparse Mixture of Experts as Unified Competitive Learning

Abstract
Sparse Mixture of Experts (SMoE) improves the efficiency of large language model training by directing input tokens to a subset of experts. Despite its success in generation tasks, its generalization ability remains an open question. In this paper, we demonstrate that current SMoEs, which fall into two categories: (1) Token Choice ;and (2) Expert Choice, struggle with tasks such as the Massive Text Embedding Benchmark (MTEB). By analyzing their mechanism through the lens of competitive learning, our study finds that the Token Choice approach may overly focus on irrelevant experts, while the Expert Choice approach risks discarding important tokens, potentially affecting performance. Motivated by this analysis, we propose Unified Competitive Learning SMoE (USMoE), a novel and efficient framework designed to improve the performance of existing SMoEs in both scenarios: with and without training. Extensive experiments across various tasks show that USMoE achieves up to a 10% improvement over traditional approaches or reduces computational inference costs by 14% while maintaining strong performance.
Sparse Mixture of Experts as Unified Competitive Learning
Giang Dothanks: Corresponding author  Hung Le  Truyen Tran Applied Artificial Intelligence Institute (A2I2), Deakin University {s224363215,thai.le,truyen.tran}@deakin.edu.au
1 Introduction
Sparse Mixture of Experts (SMoE) models have achieved notable success in natural language processing (NLP) and visual representation learning tasks (Du et al., 2022; Fedus et al., 2022; Riquelme et al., 2021a; Shen et al., 2023). These advancements build on the Transformer architecture (Vaswani et al., 2017) and its variants (Child et al., 2019; Dai et al., 2019b), which leverage large datasets and significant compute resources. However, training large Transformer models can be prohibitively expensive, requiring extensive compute hours (Kaddour et al., 2023). To overcome this issue, SMoE models activate only a subset of experts for each input, reducing inference time compared to dense models (Shazeer et al., 2017a; Zoph et al., 2022; Artetxe et al., 2022; Krajewski et al., 2024). The SMoE architecture can be categorized into two variants: Token Choice, which assigns experts to each token Dai et al. (2024); Team (2024); Muennighoff et al. (2024); Jiang et al. (2024a), and Expert Choice, which assigns tokens to each expert Zhou et al. (2022b). The advantage of Token Choice lies in its ability to dynamically select experts for each token, while Expert Choice ensures a more balanced token distribution across experts.
Figure 1: We compare the performance of USMoE (ours) with the Expert Choice and Token Choice approaches on the Massive Text Embedding Benchmark (MTEB). The results demonstrate that our method outperforms traditional approaches across six tasks using OLMoE-7B  Muennighoff et al. (2024) without additional training.
Despite their promising results, SMoE models have several limitations. The Expert Choice approach suffers from token dropping  Zhou et al. (2022b), while the Token Choice approach struggles with unbalanced expert loading  Shazeer et al. (2017b). Additionally, both approaches are prone to representation collapse, where either a few experts dominate the routing or all experts learn similar representations Chi et al. (2022a); Chen et al. (2022). Recent research has explored improving router policies Chi et al. (2022b); Chen et al. (2023a); Do et al. (2023a) to mitigate these issues. However, existing methods face two key challenges: (1) the use of auxiliary losses requires balancing router loss and task loss, leading to trade-offs, and (2) they still struggle with the fundamental limitations of either the Token Choice or Expert Choice approach.
Which approach is better for SMoE - Token Choice or Expert Choice - in terms of generalization across multiple tasks, both with and without training?
In this paper, we address this question by reexamining SMoE through the lens of Competitive Learning Rumelhart and Zipser (1985); Kaski and Kohonen (1994); Srivastava et al. (2013); Pham et al. (2024a). From this perspective, Token Choice can be seen as horizontal competitive learning, where the most similar expert is selected for each token, while Expert Choice represents vertical competitive learning, where each expert selects the most similar tokens. This viewpoint reveals a key trade-off: Expert Choice risks dropping important tokens, whereas Token Choice must process both relevant and irrelevant tokens.
Building on this analysis, we propose Unified Competitive Learning SMoE (USMoE), a robust and efficient framework comprising two key components: (1) Unified Competitive Score and (2) Unified Competitive Mechanism. These components enable the SMoE model to dynamically prioritize tokens or experts while ensuring the selection of the most similar token-expert pair, enhancing both robustness and effectiveness. To demonstrate the effectiveness of our approach, we evaluate USMoE across multiple scenarios, including pretraining and both fine-tuned and non-fine-tuned settings. USMoE consistently outperforms baseline methods across these scenarios, with particularly strong gains in tasks that require deep input understanding, such as semantic textual similarity, classification, and clustering. Extensive experiments across various benchmarks show that USMoE achieves up to a 10% improvement over traditional approaches or reduces inference computational costs by 14%, all while maintaining high performance.
In summary, this paper makes the following key contributions:
  • We introduce a Competitive Learning perspective on SMoE, highlighting the weaknesses of existing approaches.
  • We propose USMoE, a robust and efficient framework that addresses the limitations of both Token Choice and Expert Choice.
  • We theoretically demonstrate that USMoE effectively mitigates representation collapse, outperforming baseline methods.
  • We conduct extensive experiments on large language models, covering pretraining and both fine-tuned and non-fine-tuned settings, providing a detailed analysis of USMoE’s performance and effectiveness.
2 Related Work
Sparse Mixture of Experts (SMoE). Sparse Mixture of Experts (SMoE), an extension of the Mixture of Experts framework (Jacobs et al., 1991; Jordan and Jacobs, 1994), has gained traction with large language models and has since been applied in various domains, including computer vision and speech recognition (Zhou et al., 2022c; Riquelme et al., 2021b). The SMoE architecture consists of two main variants: Token Choice, where experts are assigned to each token (Shazeer et al., 2017b; Fedus et al., 2022; Jiang et al., 2024b; Do et al., 2024b), and Expert Choice, where tokens are assigned to specific experts (Zhou et al., 2022b).
Token Choice treats all tokens equally, which has raised concerns among researchers (Wu et al., 2021; Hou et al., 2022; Lin et al., 2025), while Expert Choice suffers from token-dropping issues. Additionally, SMoE faces the challenge of representation collapse, where experts produce similar outputs. Various solutions have been proposed, such as XMoE, which employs low-dimensional routing scores (Chi et al., 2022b), and SMoE-dropout, which gradually activates more experts (Chen et al., 2023a). Other approaches, including HyperRouter (Do et al., 2023a) and StableMoE (Dai et al., 2022a), focus on enhancing router stability and robustness. Although these advancements have improved SMoE models, representation collapse remains a persistent issue (Pham et al., 2024a; Do et al., 2024a). Our approach addresses this by optimizing the alignment between tokens and the most suitable experts, expanding expert specialization and mitigating collapse.
Competitive Learning. Competitive learning is an unsupervised learning approach where computational units compete to respond to a given input, enabling feature discovery (McClelland and Rumelhart, 1987; Andersen et al., 1969; Stefanis and Jasper, 1969). Inspired by the biological brain, this concept has recently been leveraged to enhance the efficiency of Large Language Models (LLMs) (Cai et al., 2024; Zhao et al., 2024). Furthermore, competitive learning has been shown to enhance the efficiency of Mixture of Experts (MoE) models (Ahn and Sentis, 2021; Cai et al., 2023).
The most relevant work to our work is CompeteSMoE (Pham et al., 2024b), which frames competition through expert outputs and employs a router trained to predict competition outcomes in a scheduled manner. However, it shares a key limitation with traditional competitive learning frameworks, activating all experts, which becomes infeasible for large-scale models with millions of parameters. In contrast, we redefine competition among expert embeddings into two categories: Token Choice SMoE, which represents horizontal competitive learning, and Expert Choice SMoE, which represents vertical competitive learning. Building on this perspective, we introduce two novel scoring methods and mechanisms - Unified Competitive Score and Unified Competitive Mechanism - that enhance traditional SMoE in both performance and efficiency.
3 Methodology
We introduce Unified Competitive Learning SMoE (USMoE), a novel and efficient Sparse Mixture of Experts framework designed to address the limitations of both Token Choice and Expert Choice through a unified competitive learning mechanism. As shown in Section 3.2, our approach consists of two key components:
  • Unified Competitive Score – a scoring function that balances expert selection.
  • Unified Competitive Mechanism – a structured routing strategy that ensures efficient and effective expert allocation.
3.1 Preliminaries
Sparse Mixture of Experts.  The Sparse Mixture of Experts (SMoE) is typically a transformer architecture that replaces the multi-layer perceptron (MLP) layers in standard transformers with Mixture of Experts (MoE) layers, inspired by (Shazeer et al., 2017a). Given $\mathbf{x}\in\mathbb{R}^{b\times d}$ as the output of the multi-head attention (MHA) layer, the result of the SMoE with $n$ experts is a weighted sum of each expert’s computation, $E_{i}(x)$, weighted by the router function $\mathcal{S}(x)$:
$f_{\mathrm{SMoE}}(\boldsymbol{x})=\sum_{i=1}^{n}\mathcal{S}(\boldsymbol{x})_{i}\cdot E_{i}(\boldsymbol{x})$ (1)
Where $\mathcal{S}(x)$ is computed by $TopK$ function as below the Equation  2, and $W_{e}$ is a learnable experts embeddings.
$\mathcal{S}(\boldsymbol{x})=\operatorname{TopK}(\operatorname{softmax}(W_{e}\cdot x),k)$ (2)
$\operatorname{TopK}(\boldsymbol{v},k)=\begin{cases}\boldsymbol{v_{i}}&\text{if}\boldsymbol{v_{i}}\text{ is in the top }k\text{ largest of }\boldsymbol{v},\\ -\infty&\text{otherwise}.\end{cases}$
3.2 Unified Competitive Learning SMoE (USMoE)
Unified Competitive Score.  As illustrated in Figure 2, the Token Choice approach selects and promotes the best expert by comparing the similarity scores between a token and the available experts. Given a representation $\mathbf{x}\in\mathbb{R}^{b\times l\times d}$, which is the output of the multi-head attention (MHA) layer, and a router matrix $W_{e}\in\mathbb{R}^{d\times n}$, where $n$ represents the number of experts, the similarity score $s_{t}$ in the Token Choice approach is computed as:
$s_{t}=\operatorname{softmax}(W_{e}\cdot x,d=-1)$ (3)
where $d$ denotes the dimension along which the softmax function is applied. In contrast, the Expert Choice approach selects and promotes the best tokens by comparing the similarity scores between an expert and the available tokens. The similarity score $s_{e}$ in the Expert Choice approach is computed as:
$s_{e}=\operatorname{softmax}(W_{e}\cdot x,d=1)$ (4)
While Expert Choice ensures an equal distribution of tokens across experts, Token Choice guarantees that each token is fairly processed by the model. To leverage the strengths of both approaches, we introduce the Unified Competitive Score, a soft combination of the Token Choice and Expert Choice scores, defined as:
$s_{u}=\alpha\cdot s_{e}+(1-\alpha)\cdot s_{t}$ (5)
where $\alpha\in[0,1]$ is a tunable control parameter that can be adjusted based on the data. In practice, we find that $\alpha\approx 0.5$ is an appropriate choice.
Unified Competitive Mechanism.  Token Choice selects experts using horizontal competitive learning, where each token is assigned to the most similar expert. This approach performs well when there is a highly relevant expert for a given token. However, its effectiveness diminishes when the similarity scores between the token and all experts are low. In contrast, Expert Choice selects tokens using vertical competitive learning, where each expert is assigned the most similar tokens. This approach is effective when the similarity score distribution across experts is well-structured. However, this condition is difficult to achieve due to the issue of representation collapse  Chi et al. (2022b). As a result, Expert Choice may either discard important tokens or lead to multiple experts selecting the same token. USMoE addresses these challenges by introducing the Unified Competitive Mechanism, which treats expert and token selection as a joint (expert, token) pairing process. As demonstrated in Algorithm  1, we first flatten the similarity matrix and then select the top-$N$ highest-scoring pairs, representing the most similar expert-token combinations.
Let the ordered array $(S^{t}_{1},S^{t}_{2},\dots,S^{t}_{N})$ represent the top-$N$ token scores selected using the Token Choice approach, where $S^{t}_{1}\leq S^{t}_{2}\leq\dots\leq S^{t}_{N}$. Similarly, let $(S^{e}_{1},S^{e}_{2},\dots,S^{e}_{N})$ denote the top-$N$ token scores selected using the Expert Choice approach, where $S^{e}_{1}\leq S^{e}_{2}\leq\dots\leq S^{e}_{N}$. Finally, let $(S^{u}_{1},S^{u}_{2},\dots,S^{u}_{N})$ represent the top-$N$ token scores selected using the Unified Competitive Score approach, where $S^{u}_{1}\leq S^{u}_{2}\leq\dots\leq S^{u}_{n}$. We derive the following inequalities, with a detailed proof provided in the Appendix A.1:
$S^{t}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N]$ (6)
$S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N]$
Based on Inequality 6, the Unified Competitive Mechanism ensures that the selected expert-token pairs are at least as optimal as those chosen by the Token Choice approach. Moreover, this mechanism addresses the limitations of Expert Choice, which may drop important tokens, and Token Choice, which may select irrelevant tokens.
Figure 2: An illustration of our USMoE framework, which enhances existing SMoE models through a Unified Competitive Learning approach. The method first flattens the score matching between tokens and experts into a 1D representation, then selects the Top-N best pairs, and finally maps the results back to the original 2D matching scores. Best viewed in color.
Algorithm 1 USMoE Layer
1:$X\in\mathbb{R}^{B\times L\times D}$, router weights $R\in\mathbb{R}^{D\times N}$, experts, controlling factor $\alpha$
2:Output $Y$
3:Compute logits:
4:      $\text{logits}\leftarrow X\cdot R$triangleright Dot product Similarity
5:Compute token choice score:
[…]
7:Compute expert choice score:
[…]
9:Compute unified score:
[…]
11:Reshape score:
12:      $U\leftarrow\operatorname{reshape}(U,B,-1)$
13:Compute top-$N$ indices and values:
[…]
15:      $\operatorname{TopK}(U,\text{topn},\text{dim}=1)$
16:Compute output using SMoE:
17:      $Y\leftarrow$
[…]
19:return $Y$
3.3 Theoretical Analysis for Representation Collapse of Sparse Mixture of Experts
Following  (Chi et al., 2022a) and  (Do et al., 2023b), we illustrate the representation collapse issue using the Jacobian matrix approach. Specifically, the Jacobian matrix of the SMoE with respect to $x\in\mathbb{R}^{b\times d}$ is given as:
$\boldsymbol{J_{SMoE}}=\mathcal{S}(x)_{k}\boldsymbol{J}^{\mathrm{FFN}}+\sum_{j=1}^{n}\mathcal{S}(x)_{k}\left(\delta_{kj}-S_{j}\right)\boldsymbol{E}(x)_{i}\boldsymbol{e}_{j}^{\top}$
$\Longrightarrow\boldsymbol{J_{SMoE}}=\mathcal{S}(x)_{k}\boldsymbol{J}^{\mathrm{FFN}}+\sum_{j=1}^{n}\boldsymbol{c}_{j}\boldsymbol{e}_{j}^{\top},$ (7)
where $\boldsymbol{c}_{j}=\mathcal{S}(x)_{k}\left(\delta_{kj}-S_{j}\right)\boldsymbol{E}(x)_{i}$. The first part of Equation  7, $\mathcal{S}(x)_{k}\boldsymbol{J}^{\mathrm{FFN}}$, represents the contribution from the input token and experts to the final output. The second part, (2) $\sum_{j=1}^{n}\boldsymbol{c}_{j}\boldsymbol{e}_{j}^{\top}$ relates to learning an improved gating function to minimize task loss. Furthermore, Equation  7 should be updated as a linear combination of expert embeddings. Due to $n<<d$ in practice, the above equation illustrates the representation.
Given $\mathcal{S}(x)=\alpha\times\mathcal{S}_{e}(x)+(1-\alpha)\times\mathcal{S}_{t}(x)$, where $\mathcal{S}_{e}(x)$ and $\mathcal{S}_{t}(x)$ represent the similarity score functions for expert choice and token choice, respectively, the Jacobian matrix of USMoE with respect to $x\in\mathbb{R}^{b\times d}$ is expressed as:
[…]
$\Longrightarrow\boldsymbol{J_{U}}=J_{1}+\sum_{j=1}^{n}c_{j}e_{j}^{\top}+\sum_{j=1}^{n}d_{j}e_{j}^{\top}$ (8)
where
$J_{1}=\mathcal{S}(x)_{k}\boldsymbol{J}^{\mathrm{FFN}},$ (9)
$c_{j}=\mathcal{S}_{e}(x)_{k}\left(\delta_{kj}-\mathcal{S}_{e}(x)_{j}\right)\boldsymbol{E}(x)_{k}\boldsymbol{e}_{j}^{\top},$ (10)
$d_{j}=\mathcal{S}_{t}(x)_{k}\left(\delta_{kj}-\mathcal{S}_{t}(x)_{j}\right)\boldsymbol{E}(x)_{k}.$ (11)
Similar to the Jacobian matrix of SMoE as Equation  7, the Jacobian matrix of USMoE also consists of two terms: (1) $J_{1}$, which depends on the input token and experts for the final output; and (2) $\sum_{j=1}^{2n}o_{j}\boldsymbol{e}_{j}^{\top}$ indicates to learn better gating function to minimize the task loss. Since $2n>>n$, USMoE is more effective than SMoE in addressing the representation collapse issue.
4 Experiments
We evaluate our method in two scenarios: (1) without additional training and (2) with training. For the first scenario, inspired by  Li and Zhou (2024), we test our method as a plug-in framework on well-trained SMoE models, including OLMoE-1B-7B Muennighoff et al. (2024):, DeepSeekMoE-16B  Dai et al. (2024), Qwen1.5-MoE-A2.7B Team (2024). We evaluate performance on a subset of tasks from the Massive Text Embedding Benchmark (MTEB)  Muennighoff et al. (2023), which covers key downstream applications for sentence embeddings, including Classification, Clustering, Pair Classification, Re-ranking, Retrieval, Semantic Textual Similarity (STS), and Summarization. Following the MTEB evaluation framework, we use Accuracy for Classification, V-Measure for Clustering, Average Precision for Pair Classification, Mean Average Precision for Re-ranking, nDCG for Retrieval, and Spearman’s correlation for STS and Summarization.
For the second scenario, we perform language model pre-training on diverse datasets, including Enwik8, Text8 Mahoney (2011), Wikitext-103 Merity et al. (2017), and One Billion Words Chelba et al. (2014). To assess performance, we fine-tune the pre-trained models on various downstream benchmarks.
4.1 Experiment Setting
We evaluate our method against two conventional SMoE approaches: (1) the Token Choice approach and (2) the Expert Choice approach, under three different settings: (1) without additional training, (2) pre-training, and (3) fine-tuning.
Without Training.  We evaluate our approach on three state-of-the-art Sparse Mixture of Experts (MoE) models: (1) OLMoE-1B-7B Muennighoff et al. (2024), which has 7B parameters, 16 layers, and 64 experts per layer; (2) Qwen1.5-MoE-A2.7B Team (2024), comprising 7B parameters, 24 layers, and 60 experts per layer; and DeepSeekMoE-16B which consists of 16B parameters, 28 layers, and 64 experts per layer. We evaluate these models on the Massive Text Embedding Benchmark (MTEB) without additional fine-tuning. Additionally, we compare methods both with and without prompts, including PromptEOL Jiang et al. (2024c). Inspired by  Li and Zhou (2024), we also compare our method with an approach that uses the similarity score between router and expert embeddings as the hidden representation, which we refer to as "Router Embedding" or simply "Router" Additionally, we evaluate against MoEE  Li and Zhou (2024), which leverages both Router Embedding and the hidden representation of the SMoE model as embeddings.
Pre-training.  To assess the effectiveness of our method, we compare USMoE with the Token Choice approaches, including SMoE Jiang et al. (2024b), SMoE-Dropout (abbreviated as "SMoE-DR"), XMoE Chi et al. (2022b), and StableMoE Dai et al. (2022a), as well as the Expert Choice approach Zhou et al. (2022a) for pre-training and fine-tuning tasks. We follow the approach of Chen et al. (2023b) and use a base Transformer-XL Dai et al. (2019a) with four decoder layers. We train both base and large-scale versions of Transformer-XL on four datasets (Enwik8, Text8, Wikitext-103, and One Billion Words) for 100k iterations, following the implementation in Chen et al. (2023b).
Fine-tuning.  We fine-tune the pre-trained weights for text classification tasks, including SST-2 Socher et al. (2013), SST-5 Socher et al. (2013), IMDB Maas et al. (2011), and BANKING77 Casanueva et al. (2020). More implementation details and additional results are provided in the Appendix A.
4.2 Without Training
Our method consistently demonstrates performance improvements across a range of MTEB tasks in two scenarios: (1) with PromptEOL Jiang et al. (2023) (denoted as "with prompts" for brevity), as shown in Table 1, and (2) without prompts, as shown in Table 2. Detailed results for datasets under each task type are provided in Appendix A. USMoE outperforms both the Expert Choice and Token Choice approaches in most cases, underscoring the complementary nature of these two methods.
Figure 1 illustrates the effectiveness of our method across various MTEB tasks using the OLMoE-1B-7B model, in comparison to both the Expert Choice and Token Choice approaches. For tasks evaluated with prompts, USMoE achieves the highest average performance across models, showing notable improvements of 4.2%, 9.1%, and 6.6% for OLMoE-1B-7B, Qwen1.5-MoE-A2.7B, and DeepSeekMoE-16B, respectively, without requiring any additional training, as detailed in Table 1. Notably, DeepSeekMoE-16B demonstrates a significant improvement from 50.1% (Token Choice) to 64.3% (USMoE), reflecting a 14.2% gain in the STS task.
For tasks evaluated without prompts, USMoE proves even more effective at enhancing the Token Choice approach, delivering notable gains of 9.3%, 7.0%, and 8.0% for OLMoE-1B-7B, Qwen1.5-MoE-A2.7B, and DeepSeekMoE-16B, respectively, across MTEB tasks without additional training. Specifically, OLMoE-1B-7B achieves a remarkable improvement from 24.1% (Token Choice) to 47.5% (USMoE), representing a 23.4% gain. This trend persists across Qwen1.5-MoE and OLMoE, where USMoE consistently outperforms both the Token Choice and Expert Choice approaches.
Interestingly, Router Embedding is less affected by prompting on the Classification dataset, as shown in Figure 3(a), while Token Choice, Expert Choice, and USMoE (ours) achieve significant performance improvements in the prompting setting. Figure 3(a) also demonstrates that our method is not only more effective but also more stable than the baselines, as indicated by a lower standard deviation. Additionally, Figure 3(b) illustrates the distribution of our method across MTEB tasks in both prompted and non-prompted scenarios. Overall, our approach outperforms the baselines in terms of performance while exhibiting lower variance across multiple tasks and different runs.
Model Task Router TC EC MoEE USMoE
OLMoE-1B-7B Classification 43.1 57.7 56.2 51.7 59.2
Clustering 16.2 24.8 26.9 23.2 30.5
PairClassification 53.5 62.0 58.9 66.0 66.8
Reranking 41.7 51.3 51.0 53.2 54.7
STS 49.4 63.5 44.2 67.8 71.1
Summarization 25.6 28.9 29.7 30.4 30.9
Average 38.3 48.0 44.5 48.7 52.2
Qwen1.5-MoE-A2.7B Classification 48.8 58.0 35.2 54.0 59.2
Clustering 14.3 34.2 29.2 30.1 35.8
PairClassification 51.9 60.5 56.0 60.3 63.4
Reranking 41.0 46.6 45.0 51.1 53.7
STS 48.3 50.1 50.0 64.3 66.3
Summarization 27.0 23.0 21.9 27.3 48.4
Average 38.6 45.4 39.6 47.9 54.5
DeepSeekMoE-16B Classification 48.6 56.4 55.4 53.0 57.3
Clustering 17.8 29.0 20.3 28.5 31.9
PairClassification 57.4 59.8 53.8 63.3 65.3
Reranking 43.8 45.7 40.9 50.6 52.1
STS 52.8 49.0 37.1 63.4 66.0
Summarization 29.1 24.4 25.7 29.2 30.9
Average 41.6 44.0 38.9 48.0 50.6
Table 1: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across across MTEB Tasks with PromptEOL Jiang et al. (2023). The best result for each row is highlighted in bold.
Model Task Router TC EC MoEE USMoE
OLMoE-1B-7B Classification 41.2 43.4 44.9 41.8 47.0
Clustering 13.7 14.7 12.0 14.5 16.9
PairClassification 45.3 39.1 35.5 45.7 50.1
Reranking 37.5 37.4 35.3 39.5 43.1
STS 39.9 24.1 18.2 39.9 47.5
Summarization 28.4 20.9 21.1 29.8 30.8
Average 34.3 29.9 27.8 35.2 39.2
Qwen1.5-MoE-A2.7B Classification 43.8 50.3 25.5 47.7 51.2
Clustering 13.6 27.4 23.2 25.2 26.7
PairClassification 45.9 46.9 43.4 51.5 53.1
Reranking 39.6 45.3 41.6 48.5 48.2
STS 38.8 38.0 35.6 51.8 54.8
Summarization 28.3 13.4 15.1 31.2 29.7
Average 35.0 36.9 30.7 42.6 43.9
DeepSeekMoE-16B Classification 43.4 46.6 44.7 44.4 46.8
Clustering 13.4 18.1 13.5 17.8 21.9
PairClassification 45.5 40.9 37.1 46.1 51.9
Reranking 38.5 38.9 35.1 42.2 45.7
STS 37.7 26.3 23.3 40.2 45.9
Summarization 24.9 22.0 18.5 24.4 28.4
Average 33.9 32.1 28.7 35.9 40.1
Table 2: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks without prompts and models. The best result for each row is highlighted in bold.
(a) Performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) on Classification Task.
(b) Results distribution of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks
Figure 3: Illustration of comparing the performance of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB tasks and three SMoE models. Each benchmark is run 10 times, reporting both the mean and standard deviation to highlight the performance and stability of our method compared to the baselines.
Transformer-XL(20M) Enwik8 Text8 WikiText-103 lm1b
[…]
SMoE-DR 1.56 1.56 58.37 93.17
XMoE 1.21 1.28 30.34 58.33
StableMoE 1.20 1.28 29.97 58.25
[…]
Table 3: Performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) across multiple datasets, with BPC on the Enwik8 and Text8 test sets, and perplexity on the WikiText-103 and One Billion Word test sets. Lower values are better, with the best results highlighted in bold.
4.3 Pre-training Result
Base models training. Table 3 summarizes the pre-training results across four datasets (enwik8, text8, WikiText-103, and One Billion Words). USMoE consistently outperforms the expert choice baseline and token choice routing methods such as XMoE Chi et al. (2022a) and StableMoE Dai et al. (2022b) on all four datasets. The strength of USMoE lies in its inference efficiency, achieved by leveraging fewer experts. Notably, on text8, USMoE surpasses SMoE while utilizing only one and a half experts. Additionally, it outperforms SMoE-Dropout (which employs two experts) on One Billion Words, lowering perplexity from 93.17 to 57.55 with the same reduced expert count. Furthermore, with only one and a half experts, USMoE reduces FLOPs by 14% compared to SMoE and SMoE-Dropout, which rely on two experts, all while maintaining strong performance.
Large models training.  USMoE not only delivers strong performance in base model training but also remains highly competitive at a large scale. Table 4 presents perplexity (PPL) results on the WikiText-103 and One Billion Words datasets using a large Transformer-XL model with 15 SMoE layers, 16 experts, and 420M parameters. The performance gap between USMoE and the baselines becomes even more pronounced at this scale, highlighting its strong scalability with increasing model complexity. Regardless of backbone size or the number of activated experts, USMoE consistently outperforms all baselines, demonstrating its effectiveness in scaling up large language models.
Transformer-XL(420M) WikiText-103 lm1b
[…]
1 31.70 35.52 25.48 58.65 65.43 56.90
2 22.42 23.30 22.06 44.56 43.39 40.53
4 23.57 23.60 22.65 45.52 43.70 40.90
8 24.20 24.37 22.88 46.36 44.22 43.24
Table 4: Large Scale performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) across multiple datasets, with perplexity on the WikiText-103 and One Billion Word test sets. Lower values are better, with the best results highlighted in bold.
4.4 Fine-tuning Result
[…]
SMoE-DR 78.6 34.4 83.5 66.7
XMoE 76.7 35.3 83.3 67.4
StableMoE 77.7 34.3 83.9 60.8
[…]
Table 5: Accuracy performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) after fine-tuned on various datasets. Higher is better, best results are in bold.
Fine-tuning.  We report the results of the fine-tuning experiment on the SST-2, SST-5, IMDB, and BANKING77 datasets in Table  5, using Transformer-XL pre-trained on enwik8. Overall, USMoE consistently achieves higher accuracy compared to other baselines across all datasets. The results demonstrate that our method is not only effective for pre-training tasks but also performs effectively on existing pre-trained models.
4.5 Ablation Studies
We investigate the effectiveness and robustness of USMoE to the different hyper-parameter settings.
4.5.1 Competitive Learning Strategy Comparison
Model Dataset TC EC USMoE-Sequence USMoE-Batch
DeepSeekMoE-16B Emotion 27.4 26.5 27.8 27.4
Toxic 60.4 58.1 60.1 59.2
Tweet 51.9 49.5 52.5 51.7
Table 6: Competitive Learning Strategy comparison of USMoE, Token Choice (TC), and Expert Choice (EC) on the classification task. Higher values are better, with the best results highlighted in bold.
The Unified Competitive Mechanism is implemented using two approaches: (1) a sequence-based method that compares all tokens within a sequence (referred to as "USMoE-Sequence") and (2) a batch-based method that compares all tokens within a batch or mini-batch (referred to as "USMoE-Batch"). We evaluate both approaches on the Classification task, with results presented in Table 6. The findings indicate that both methods outperform the Expert Choice and Token Choice approaches, demonstrating the effectiveness of our method. Notably, the sequence-based approach achieves superior performance in the Classification task, as it ensures that no important tokens are missed within a sequence - an assurance that the batch/mini-batch implementation may not always provide.
4.5.2 Robustness to the controlling factor
Model Dataset $\alpha$
0.0 0.3 0.5 0.7 0.9 1.0
DeepSeekMoE-16B Emotion 27.4 27.1 27.8 27.6 27.7 26.5
Toxic 60.4 60.0 60.1 56.8 57.3 58.1
Tweet 51.9 53.2 52.5 53.3 52.9 49.5
Table 7: Performance comparison of DeepSeekMoE-16B across different classification datasets with varying $\alpha$ values. Higher is better; best results are in bold.
The Unified Competitive Score ($\alpha$) enables the model to adjust its scoring mechanism, either favoring a diverse set of experts per sequence or distributing experts more evenly across tokens. We evaluate the robustness of the controlling factor $\alpha$ on the classification task using the DeepSeekMoE-16B model, with results presented in Table 7. When $\alpha=0.0$, the scoring mechanism aligns with Token Choice, while at $\alpha=1.0$, it follows Expert Choice.
Overall, USMoE demonstrates strong effectiveness within the range of $\alpha\in(0.3,0.7)$, striking a balance between expert diversity and token importance. This range provides an optimal trade-off between enforcing SMoE’s competition policy and enhancing traditional approaches for the task, leading to superior overall performance. Notably, all tested $\alpha$ configurations outperform the Expert Choice approach ($\alpha=1.0$).
Summary and hyper-parameter configuration guidelines. Our experiments demonstrate that USMoE not only excels in pre-training and transfer learning but also achieves superior generalization compared to traditional approaches, all without requiring additional training. Ablation studies in Section 4.5 indicate that USMoE is highly effective within the range of $\alpha\in(0.3,0.7)$. This finding provides a practical guideline for efficiently tuning USMoE through cross-validation.
5 Conclusion
In this research, we reformulate Token Choice and Expert Choice Sparse Mixture of Experts from a competitive learning perspective, highlighting their limitations. Building on this analysis, we introduce Unified Competitive Learning SMoE (USMoE) - a novel and efficient framework that enhances SMoE through a unified competitive learning approach.
We theoretically prove that USMoE achieves superior expert selection compared to traditional methods, effectively improving expert learning capacity while mitigating expert collapse. As a result, USMoE learns more robust expert representations and overcomes the representation collapse issues commonly observed in conventional SMoE training.
Experiments across both training-free and training-based settings (including pre-training and fine-tuning) demonstrate that USMoE enables more efficient and effective training and inference compared to state-of-the-art routing strategies.
Limitations
Our study focuses on enhancing the efficiency and effectiveness of Large Language Models (LLMs) through SMoE. Our approach proves effective in both training and non-training settings. While the results are promising, our pre-training experiments were constrained by computational resources, limiting us to medium-scale datasets and a base Transformer-XL model. Consequently, further empirical evaluation is required to assess the scalability of USMoE beyond 100B parameters and compare it with other SMoE strategies in modern LLMs.
Ethics Statement
Despite encouraging results, training large-scale LLMs remains highly resource-intensive, requiring careful management of computational costs. Additionally, our study relies on web-sourced data, which is known to contain gender and racial biases, highlighting the need for further efforts to mitigate these issues. Lastly, while our work represents a significant step toward advancing LLMs development, it also emphasizes the importance of robust regularization to prevent potential misuse in harmful applications.
Appendix A Appendix
Supplementary Material for “Sparse Mixture of Experts as Unified Competitive Learning”
This document is organized as follows: Appendix A.1 provides the theoretical proof supporting the discussion in Section 3.2. Following this, Appendix A.3 offers a detailed analysis of the SMoE router and explains why our method outperforms the baseline approaches. Appendix A.2 presents supplementary experimental results, and Appendix A.4 describes the implementation details in full.
A.1 Theoretical Proof for Section  3.2
Let $x\in\mathbb{R}^{N\times d}$ represent an input token embedding and $\mathcal{E}\in\mathbb{R}^{d\times n}$ represent the expert embeddings, where $d$ is the dimension of the SMoE model and $n$ is the number of experts. The similarity score $S$ between the input token and the expert embeddings is computed using a dot product:
$S=x\cdot\mathcal{E}.$ (12)
Since the softmax function is monotonic, the selection of top-$k$ values remains unchanged under softmax transformation. Formally, we have:
$\operatorname{TopK}(\operatorname{softmax}(S),k)=\operatorname{TopK}(S,k).$ (13)
By setting $k=1$, the top similarity score for a given token using the Token Choice approach is:
$S^{t}_{j}=\max(S_{ik}),\quad\forall k\in[1,n].$ (14)
Since this operation is performed independently for each input token $i$, we extend this to all tokens $q$ in the batch:
$S^{t}_{j}\leq\max(S_{qk}),\quad\forall k\in[1,n],\quad\forall q\in[1,N].$ (15)
By definition, the Unified Competitive Mechanism approach selects the highest similarity score across all tokens and experts:
$S^{u}_{j}=\max(S_{qk}),\quad\forall k\in[1,n],\quad\forall q\in[1,N].$ (16)
Thus, we establish the first inequality:
$S^{t}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N].$ (17)
Similarly, in the Expert Choice approach, each expert selects the best matching token, leading to:
$S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N].$ (18)
Therefore, we conclude:
$S^{t}_{i}\leq S^{u}_{i},\quad S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N].$ (19)
This completes the proof. $\square$
A.2 Additional Results
We provide a detailed evaluation of three state-of-the-art SMoE models: OLMoE-1B-7B (Table  8), Qwen1.5-MoE-A2.7B (Table 9), and DeepSeekMoE-16B (Table  10). Our results demonstrate the effectiveness of our method across various models and prompts, comparing its performance against baseline approaches such as Token Choice (TC) and Expert Choice (EC).
Category Model Dataset Setting Router TC EC MoEE USMoE
Classification OLMoE-1B-7B Emotion None 24.1 24.5 26.3 25.1 28.2
Prompt 27.6 49.9 49.0 44.5 51.7
Toxic None 51.9 58.9 59.9 51.9 60.6
Prompt 52.3 65.2 61.3 53.4 66.6
Tweet None 47.7 46.8 48.3 48.4 52.1
Prompt 49.5 58.0 58.4 57.2 59.2
Clustering OLMoE-1B-7B Medrxiv None 15.0 17.6 14.8 17.4 21.2
Prompt 15.8 23.9 27.7 22.0 26.3
20Groups None 12.4 11.8 9.2 11.5 12.6
Prompt 16.7 25.7 26.2 24.4 34.6
Pair Classification OLMoE-1B-7B SemEval None 43.6 35.8 31.3 43.6 45.7
Prompt 45.7 46.7 40.9 53.8 54.2
URLCorpus None 47.0 42.4 39.7 47.8 54.5
Prompt 61.4 77.4 76.9 78.2 79.5
Reranking OLMoE-1B-7B Ask None 41.3 41.0 39.0 41.4 42.4
Prompt 43.4 51.9 49.9 50.2 51.8
SciDocs None 45.5 46.3 46.9 50.8 59.9
Prompt 53.6 69.6 73.1 75.1 76.6
StackOver None 25.8 24.8 20.1 26.4 27.0
Prompt 28.1 32.5 30.0 34.3 35.6
STS OLMoE-1B-7B Biosses None 39.3 13.6 7.7 29.7 45.4
Prompt 51.2 61.8 67.6 70.2 74.3
SickR None 50.3 46.3 26.4 53.0 55.1
Prompt 51.9 65.7 37.6 66.1 67.9
STS12 None 40.1 8.6 11.1 37.8 44.6
Prompt 51.3 53.8 37.5 63.6 67.1
STS13 None 40.5 21.1 18.2 43.4 49.8
Prompt 52.5 66.5 40.4 72.7 76.2
STS14 None 29.5 13.4 13.3 31.7 38.7
Prompt 41.1 56.8 33.9 64.2 67.3
STS15 None 30.8 27.8 22.5 33.3 43.3
Prompt 46.4 69.3 38.4 66.4 71.7
STS16 None 46.5 38.9 28.9 45.8 51.7
Prompt 52.4 70.1 49.4 68.3 72.7
STSBen None 42.2 23.4 17.5 44.5 51.1
Prompt 48.6 63.6 48.9 70.7 71.4
Summarization OLMoE-1B-7B Medrxiv None 28.4 20.9 21.1 29.8 30.8
Prompt 25.6 28.9 29.7 30.4 30.9
Table 8: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks with OLMoE-1B-7B models. The best result for each row is highlighted in bold.
Category Model Dataset Setting Router TC EC MoEE USMoE
Classification Qwen1.5-MoE-A2.7B Emotion None 27.2 33.9 30.1 34.3 35.7
Prompt 37.0 48.5 47.4 47.2 49.9
Toxic None 53.0 61.1 21.3 52.9 61.4
Prompt 53.4 64.5 19.8 54.1 66.2
Tweet None 51.1 55.9 25.1 55.9 56.5
Prompt 56.1 61.1 38.6 60.7 61.5
Clustering Qwen1.5-MoE-A2.7B Medrxiv None 15.3 23.3 21.3 23.0 23.4
Prompt 14.2 24.6 19.8 21.8 25.1
20Groups None 12.0 31.5 25.1 27.4 30.0
Prompt 14.4 43.8 38.6 38.4 46.4
Pair Classification Qwen1.5-MoE-A2.7B SemEval None 42.0 38.8 34.7 42.5 44.2
Prompt 47.0 52.4 46.3 52.4 54.1
URLCorpus None 49.8 54.9 52.1 60.6 62.1
Prompt 56.7 68.7 65.8 68.2 72.7
Reranking Qwen1.5-MoE-A2.7B Ask None 43.1 45.8 43.4 47.3 47.5
Prompt 43.3 48.3 49.1 49.5 51.4
SciDocs None 49.6 60.6 55.3 67.0 65.6
Prompt 50.9 60.1 55.8 68.7 73.0
StackOver None 26.2 29.5 26.2 31.1 31.4
Prompt 28.8 31.3 30.2 35.2 36.6
STS Qwen1.5-MoE-A2.7B Biosses None 33.8 32.5 34.7 49.6 52.6
Prompt 55.1 55.8 48.5 68.4 66.2
SickR None 51.0 55.5 40.4 61.0 63.6
Prompt 50.2 59.7 51.1 64.3 66.3
STS12 None 40.2 16.9 18.6 46.3 48.3
Prompt 49.3 25.0 31.8 59.2 61.4
STS13 None 38.1 42.9 44.2 56.7 61.8
Prompt 53.3 57.5 54.6 73.4 75.7
STS14 None 28.1 26.5 25.6 45.4 48.9
Prompt 40.4 38.8 40.7 60.0 62.7
STS15 None 34.8 40.5 38.4 46.1 48.0
Prompt 40.7 52.3 54.2 58.8 62.5
STS16 None 47.6 51.0 48.1 58.1 59.6
Prompt 51.6 64.2 65.1 65.7 68.2
STSBen None 37.0 37.7 34.7 50.9 55.5
Prompt 45.6 47.8 54.5 64.5 67.3
Summarization Qwen1.5-MoE-A2.7B Medrxiv None 28.3 13.4 15.1 31.2 29.7
Prompt 27.0 23.0 21.9 27.3 48.4
Table 9: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks with Qwen1.5-MoE-A2.7B models. The best result for each row is highlighted in bold.
Category Model Dataset Setting Router TC EC MoEE USMoE
Classification DeepSeekMoE-16B Emotion None 26.1 27.4 26.5 27.6 27.8
Prompt 37.9 48.3 46.1 46.4 48.8
Toxic None 53.3 60.4 58.1 53.1 60.1
Prompt 53.1 62.4 62.5 53.6 64.2
Tweet None 51.0 51.9 49.5 52.6 52.5
Prompt 54.9 58.4 57.5 58.9 58.9
Clustering DeepSeekMoE-16B Medrxiv None 15.1 23.0 17.3 22.0 25.0
Prompt 17.0 25.7 20.9 24.0 25.8
20Groups None 11.7 13.2 9.7 13.7 18.7
Prompt 18.6 32.3 19.8 33.0 37.9
Pair Classification DeepSeekMoE-16B SemEval None 44.6 40.2 32.6 43.5 46.0
Prompt 48.4 47.2 46.6 51.3 54.6
URLCorpus None 46.4 41.7 41.6 48.6 57.9
Prompt 66.5 72.4 61.1 75.4 75.9
Reranking DeepSeekMoE-16B Ask None 41.7 41.1 40.1 42.3 45.0
Prompt 43.5 43.8 44.7 46.9 49.9
SciDocs None 48.2 50.6 44.7 57.1 61.9
Prompt 58.3 65.6 55.3 72.6 72.9
StackOver None 25.7 24.9 20.4 27.3 30.2
Prompt 29.7 27.6 22.6 32.3 33.6
STS DeepSeekMoE-16B Biosses None 29.5 31.7 27.7 26.8 35.4
Prompt 47.0 40.1 41.5 57.6 55.3
SickR None 50.4 47.4 29.4 53.1 56.7
Prompt 56.0 61.9 38.7 65.8 67.7
STS12 None 44.0 4.3 13.9 45.0 46.9
Prompt 57.8 31.0 28.4 64.0 64.2
STS13 None 36.0 28.4 27.5 41.1 49.1
Prompt 55.3 56.0 41.2 70.9 75.2
STS14 None 25.4 12.0 13.0 28.2 37.3
Prompt 44.9 41.0 31.1 58.6 63.8
STS15 None 34.8 33.9 25.6 38.7 40.7
Prompt 49.7 46.5 33.0 58.5 63.0
STS16 None 44.9 34.4 33.1 46.9 55.7
Prompt 56.7 58.0 44.2 64.5 69.7
STSBen None 36.6 18.3 15.8 42.1 45.6
Prompt 54.9 57.7 39.0 67.8 69.2
Summarization DeepSeekMoE-16B Medrxiv None 24.9 22.0 18.5 24.4 28.4
Prompt 29.1 24.4 25.7 29.2 30.9
Table 10: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks with DeepSeekMoE-16B models. The best result for each row is highlighted in bold.
A.3 In-depth Analysis
We visualize the router behavior of USMoE in Figure 6 and contrast it with the router behaviors of the Token Choice approach (Figure 4) and the Expert Choice approach (Figure 5). Notably, the router in the OLMoE-1B-7B model exhibits a strong preference for specific experts. For instance, in the Emotion Classification task, Experts 8, 30, and 58 are consistently prioritized in both the Token Choice and Expert Choice approaches. This bias limits the model’s adaptability and effectiveness for downstream tasks. USMoE tackles this challenge by introducing the Unified Competitive Mechanism, which promotes more balanced and diverse expert selections, as illustrated in Figure 6. This enhancement enables USMoE to outperform the baselines on the Emotion Classification task.
Figure 4: Token Choice Router visualization for the OLMoE-1B-7B model on the Emotion Classification task. The scores of selected experts are replaced with -10.0 (lower than the minimum score) to enhance visualization. Best viewed in color.
Figure 5: Expert Choice Router visualization for the OLMoE-1B-7B model on the Emotion Classification task. The scores of selected experts are replaced with -10.0 (lower than the minimum score) to enhance visualization. Best viewed in color.
Figure 6: USMoE Router visualization for the OLMoE-1B-7B model on the Emotion Classification task. The scores of selected experts are replaced with -10.0 (lower than the minimum score) to enhance visualization. Best viewed in color.
We track the number of unique experts utilized by the OLMoE-1B-7B model for each sequence in the Emotion Classification task as Figure 7(a). Our analysis reveals that the Expert Choice approach employs 11 out of 16 experts, indicating a lower level of specialization among experts. In contrast, both USMoE and the Token Choice approach use an average of 0.9 to 1 expert per sequence, demonstrating superior expert specialization. Furthermore, we analyze the token dropping behavior of the Expert Choice approach and observe a significant increase in dropping rates when scaling to larger datasets or models, such as pre-training the Transformer-XL Large model on the One Billion Word dataset, as shown in Figure 7(b). This increase in dropping rates may negatively impact model performance. In contrast, our method maintains a consistently low dropping rate (<0.1), demonstrating its superiority over the Expert Choice approach for scalability. Additionally, our method proves more robust than the Token Choice approach, as it effectively drops irrelevant tokens without compromising performance.
(a) Number Experts per Sequence of USMoE, TC, and EC on Emotion dataset.
(b) Token Dropping of USMoE, Token Choice (TC), Expert Choice (EC) for Pre-training on One Billion Word dataset.
Figure 7: Comparison of the number of experts per sequence for USMoE, Token Choice (TC), and Expert Choice (EC) on the Emotion dataset using the OLMoE-1B-7B model, along with a comparison of token dropping rates for USMoE, TC, and EC during pre-training on the One Billion Word dataset.
A.4 Implementation Details
For the Without Training experiments, we implement our method based on the publicly available MoEE implementation Li and Zhou (2024) [^1] https://github.com/tianyi-lab/MoE-Embedding. Due to resource constraints, we validate our method and the baselines using 4-bit quantization with a batch size of 128. For the OLMoE-1B-7B model, we conduct experiments on a single H100 GPU, while for the Qwen1.5-MoE-A2.7B and DeepSeekMoE-16B models, we utilize two H100 GPUs.
The base Transformer-XL variant (Chen et al., 2023b) comprises four Transformer decoder layers, each with an input dimension of 256. Each layer includes a self-attention mechanism with eight attention heads, followed by a Feed-forward Neural Network (FFN) that has an inner dimension of 512. The dropout ratio is set at 0.1. We divide the FFN into 16 experts, each with the same dimensions. For the larger variants, we scale the model up to twelve layers.
Our experiments are based on the publicly available SMoE-Dropout implementation (Chen et al., 2023b) [^2] https://github.com/VITA-Group/Random-MoE-as-Dropout. The pre-training experiments were conducted using a single H100 GPU, while the fine-tuning experiments were performed on a single A100 GPU. It is important to note that parallel training on multiple GPUs may produce different results.
A.4.1 Pre-training Experiments
We provide the USMoE implementation details for pre-training our Transformer-XL base and large on enwik8, text8, WikiText-103, and One Billion Word in Table 11.
Table 11: Implementation details for pre-training experiments on enwik8, text8, WikiText-103, and One Billion Word datasets.
Dataset Input length Batch size Optimizer Lr # Iterations
enwik8 512 48 Adam 2.5e-4 100k
text8 512 48 Adam 2.5e-4 100k
WikiText-103 512 22 Adam 2.5e-4 100k
One Billion Word 512 11 Adam 2.5e-4 100k
A.4.2 Fine-tuning Experiments
To perform the fine-tuning experiments, we utilize the same model architecture as in the pre-training phase. Table 12 presents the implementation details for the fine-tuning experiments conducted across four different datasets.
Table 12: Implementation for fine-tuning experiments on downstream tasks.
Dataset Input length Batch size Optimizer Lr # Epochs
SST-2 512 16 Adam 1e-4 15
SST-5 512 16 Adam 1e-4 15
IMDB 512 4 Adam 1e-4 15
BANKING77 512 16 Adam 1e-4 15

### v2

@@
-Sparse Mixture of Experts as Unified Competitive Learning
+Unified Sparse Mixture of Experts
 Abstract
-Sparse Mixture of Experts (SMoE) improves the efficiency of large language model training by directing input tokens to a subset of experts.
-Despite its success in generation tasks, its generalization ability remains an open question.
-In this paper, we demonstrate that current SMoEs, which fall into two categories: (1) Token Choice ;and (2) Expert Choice, struggle with tasks such as the Massive Text Embedding Benchmark (MTEB).
-By analyzing their mechanism through the lens of competitive learning, our study finds that the Token Choice approach may overly focus on irrelevant experts, while the Expert Choice approach risks discarding important tokens, potentially affecting performance.
-Motivated by this analysis, we propose Unified Competitive Learning SMoE (USMoE), a novel and efficient framework designed to improve the performance of existing SMoEs in both scenarios: with and without training.
-Extensive experiments across various tasks show that USMoE achieves up to a 10% improvement over traditional approaches or reduces computational inference costs by 14% while maintaining strong performance.
-Sparse Mixture of Experts as Unified Competitive Learning
-Giang Dothanks: Corresponding author  Hung Le  Truyen Tran Applied Artificial Intelligence Institute (A2I2), Deakin University {s224363215,thai.le,truyen.tran}@deakin.edu.au
+Sparse Mixture of Experts (SMoEs) models scale the capacity of models while maintaining constant computational overhead.
+Early designs typically relied on a fixed value of $k$, where $k$ represents either the number of experts selected per token or the number of tokens assigned per expert.
+However, these approaches encounter three key limitations: they may fail to route to important experts or tokens, may assign irrelevant ones, and often suffer from representation collapse among experts.
+This paper reexamines SMoEs through the lens of Linear Programming, and proposes a Unified Sparse Mixture of Experts (USMoE) framework that addresses these limitations.
+Specifically, our approach introduces a unified mechanism that integrates information from both the expert and token dimensions, and a unified scoring function that linearly combines similarity scores between experts and tokens.
+We provide both theoretical justification and empirical evidence demonstrating USMoE’s effectiveness in overcoming the limitations of traditional routing methods.
+Through comprehensive evaluations on both clean and corrupted settings for large language models and vision tasks, under both training-free and training scenarios, USMoE achieves up to a 10% performance improvement over standard approaches or reduces inference costs by up to 14%, while maintaining competitive accuracy.
 1 Introduction
-Sparse Mixture of Experts (SMoE) models have achieved notable success in natural language processing (NLP) and visual representation learning tasks (Du et al., 2022; Fedus et al., 2022; Riquelme et al., 2021a; Shen et al., 2023).
-These advancements build on the Transformer architecture (Vaswani et al., 2017) and its variants (Child et al., 2019; Dai et al., 2019b), which leverage large datasets and significant compute resources.
-However, training large Transformer models can be prohibitively expensive, requiring extensive compute hours (Kaddour et al., 2023).
-To overcome this issue, SMoE models activate only a subset of experts for each input, reducing inference time compared to dense models (Shazeer et al., 2017a; Zoph et al., 2022; Artetxe et al., 2022; Krajewski et al., 2024).
-The SMoE architecture can be categorized into two variants: Token Choice, which assigns experts to each token Dai et al.
-(2024); Team (2024); Muennighoff et al.
-(2024); Jiang et al.
-(2024a), and Expert Choice, which assigns tokens to each expert Zhou et al.
-(2022b).
+Sparse Mixture of Experts (SMoE) models have achieved notable success in natural language processing (NLP) and visual representation learning tasks (du_glam_2022, ; fedus_switch_2022, ; riquelme2021scalingvisionsparsemixture, ; shen-etal-2023-scaling, ).
+These advancements build on the Transformer architecture (NIPS2017_3f5ee243, ) and its variants (child2019generating, ; dai2019transformerxl, ), which leverage large datasets and significant compute resources.
+However, training large Transformer models can be prohibitively expensive, requiring extensive compute hours (kaddour2023challenges, ).
+To overcome this issue, SMoE models activate only a subset of experts for each input, reducing inference time compared to dense models (shazeer2017outrageously, ; zoph2022stmoe, ; artetxe2022efficient, ; krajewski2024scaling, ).
+The SMoE architecture can be categorized into two variants: Token Choice, which assigns experts to each token dai2024deepseekmoeultimateexpertspecialization ; qwen_moe ; muennighoff2024olmoeopenmixtureofexpertslanguage ; jiang2024mixtral , and Expert Choice, which assigns tokens to each expert zhou_mixture_experts_2022 .
 The advantage of Token Choice lies in its ability to dynamically select experts for each token, while Expert Choice ensures a more balanced token distribution across experts.
-Figure 1: We compare the performance of USMoE (ours) with the Expert Choice and Token Choice approaches on the Massive Text Embedding Benchmark (MTEB).
-The results demonstrate that our method outperforms traditional approaches across six tasks using OLMoE-7B  Muennighoff et al.
-(2024) without additional training.
+Input
+USMoE
+TC
+EC
+Input
+USMoE
+TC
+EC
+Figure 1: We compare token routing performance in vision tasks using 7×7 images, where each token is color-coded based on its assigned expert under setting $c=t$, where $c$ is a sparsity constraint and $t$ is number of image patches.
+Left: When the object is easy to distinguish, Expert Choice (EC) fails to assign different experts appropriately.
+Token Choice (TC) performs better but still does not align perfectly with the actual object, while USMoE correctly separates the object.
+Right: In more challenging images, both Expert Choice (EC) and Token Choice (TC) fail to distinguish between object and background.
+In contrast, USMoE successfully differentiates the object from the background, demonstrating greater efficiency in vision tasks compared to EC and TC, as further shown in Section 4.
 Despite their promising results, SMoE models have several limitations.
-The Expert Choice approach suffers from token dropping  Zhou et al.
-(2022b), while the Token Choice approach struggles with unbalanced expert loading  Shazeer et al.
-(2017b).
-Additionally, both approaches are prone to representation collapse, where either a few experts dominate the routing or all experts learn similar representations Chi et al.
-(2022a); Chen et al.
-(2022).
-Recent research has explored improving router policies Chi et al.
-(2022b); Chen et al.
-(2023a); Do et al.
-(2023a) to mitigate these issues.
-However, existing methods face two key challenges: (1) the use of auxiliary losses requires balancing router loss and task loss, leading to trade-offs, and (2) they still struggle with the fundamental limitations of either the Token Choice or Expert Choice approach.
-Which approach is better for SMoE - Token Choice or Expert Choice - in terms of generalization across multiple tasks, both with and without training?
-In this paper, we address this question by reexamining SMoE through the lens of Competitive Learning Rumelhart and Zipser (1985); Kaski and Kohonen (1994); Srivastava et al.
-(2013); Pham et al.
-(2024a).
-From this perspective, Token Choice can be seen as horizontal competitive learning, where the most similar expert is selected for each token, while Expert Choice represents vertical competitive learning, where each expert selects the most similar tokens.
-This viewpoint reveals a key trade-off: Expert Choice risks dropping important tokens, whereas Token Choice must process both relevant and irrelevant tokens.
-Building on this analysis, we propose Unified Competitive Learning SMoE (USMoE), a robust and efficient framework comprising two key components: (1) Unified Competitive Score and (2) Unified Competitive Mechanism.
+The Expert Choice approach suffers from token dropping  zhou_mixture_experts_2022 , while the Token Choice approach struggles with unbalanced expert loading  shazeer2017outrageouslylargeneuralnetworks .
+Additionally, both approaches are prone to representation collapse, where either a few experts dominate the routing or all experts learn similar representations chi_representation_2022 ; chen_towards_2022 .
+Recent research has explored improving router policies chi2022representation ; chen2023sparse ; do2023hyperrouter to mitigate these issues.
+However, existing methods face three key challenges: (1) The reliance on auxiliary losses requires careful balancing between the router loss and the task loss, which introduces trade-offs; (2) Token Choice (TC) struggles to handle noisy tokens effectively; and (3) Expert Choice (EC) suffers from information leakage issues, which significantly degrade performance on autoregressive models zhou2022mixtureofexpertsexpertchoicerouting ; wang2024auxiliarylossfreeloadbalancingstrategy ; raposo2024mixtureofdepthsdynamicallyallocatingcompute .
+As a result, the question of how to optimally select experts or tokens remains open.
+Research Question What are the optimal criteria for selecting experts or tokens in SMoE models under fixed computational budgets?
+Loading Balance Robustness Token Dropping Information Leak Expert Collapses
+Token Choice Φ Φ Φ
+Expert Choice Φ Φ
+USMoE (ours) Φ Φ Φ Φ Φ
+Table 1: Benefits of the proposed USMoEs’ model form over existing methods.
+In this paper, we revisit SMoEs through the lens of Linear Programming NIPS2015_09662890 ; NEURIPS2021_a8fbbd3b to better understand the criteria for expert and token selection.
+From this perspective, Token Choice selects experts along the expert dimension by assigning each token to its most similar expert, while Expert Choice selects tokens along the token dimension by allowing each expert to choose the most similar tokens.
+This perspective highlights a key trade-off: Expert Choice risks dropping important tokens, whereas Token Choice struggles to handle noisy or irrelevant tokens.
+Furthermore, both approaches suffer from the problem of representation collapse chi2022representation ; do2023hyperrouter ; pham2024competesmoe .
+Building on this analysis, we propose the Unified Sparse Mixture of Experts (USMoE), a robust and efficient framework consisting of two key components: (1) the Unified Score and (2) the Unified Mechanism.
+The Unified Score is defined as a linear combination of two mapping functions that transform the dot product similarity between token and expert embeddings into a probability distribution - using functions such as softmax or sigmoid, as shown in Algorithm 1.
+In parallel, the Unified Mechanism incorporates information from both the token and expert dimensions, as illustrated in Figure 2.
 These components enable the SMoE model to dynamically prioritize tokens or experts while ensuring the selection of the most similar token-expert pair, enhancing both robustness and effectiveness.
-To demonstrate the effectiveness of our approach, we evaluate USMoE across multiple scenarios, including pretraining and both fine-tuned and non-fine-tuned settings.
+We summarize the benefits of USMoEs’ model form over existing methods in Table  1.
+To demonstrate the effectiveness of our approach, we evaluate USMoE across a variety of scenarios, including training-free, pretraining, and fine-tuning settings, spanning both language and vision domains under both clean and corrupted datasets.
 USMoE consistently outperforms baseline methods across these scenarios, with particularly strong gains in tasks that require deep input understanding, such as semantic textual similarity, classification, and clustering.
@@
 In summary, this paper makes the following key contributions:
-  • We introduce a Competitive Learning perspective on SMoE, highlighting the weaknesses of existing approaches.
-  • We propose USMoE, a robust and efficient framework that addresses the limitations of both Token Choice and Expert Choice.
+  • We introduce a Linear Programming perspective on SMoE, highlighting the weaknesses of existing approaches.
+  • We propose USMoE, a robust and efficient framework comprising two modules: (1) the Unified Score and (2) the Unified Mechanism, which together address the limitations of conventional approaches, as summarized in Table 1.
   • We theoretically demonstrate that USMoE effectively mitigates representation collapse, outperforming baseline methods.
-  • We conduct extensive experiments on large language models, covering pretraining and both fine-tuned and non-fine-tuned settings, providing a detailed analysis of USMoE’s performance and effectiveness.
+  • We conduct extensive experiments on both large language models and vision domains, covering pretraining and both fine-tuning and training-free settings, providing a detailed analysis of USMoE’s performance and effectiveness.
 2 Related Work
 Sparse Mixture of Experts (SMoE).
-Sparse Mixture of Experts (SMoE), an extension of the Mixture of Experts framework (Jacobs et al., 1991; Jordan and Jacobs, 1994), has gained traction with large language models and has since been applied in various domains, including computer vision and speech recognition (Zhou et al., 2022c; Riquelme et al., 2021b).
-The SMoE architecture consists of two main variants: Token Choice, where experts are assigned to each token (Shazeer et al., 2017b; Fedus et al., 2022; Jiang et al., 2024b; Do et al., 2024b), and Expert Choice, where tokens are assigned to specific experts (Zhou et al., 2022b).
-Token Choice treats all tokens equally, which has raised concerns among researchers (Wu et al., 2021; Hou et al., 2022; Lin et al., 2025), while Expert Choice suffers from token-dropping issues.
+Sparse Mixture of Experts (SMoE), an extension of the Mixture of Experts framework (jacobs1991, ; jordan1994, ), has gained traction with large language models and has since been applied in various domains, including computer vision and speech recognition (NEURIPS2022_2f00ecd7, ; NEURIPS2021_48237d9f, ).
+The SMoE architecture consists of two main variants: Token Choice, where experts are assigned to each token (shazeer2017outrageouslylargeneuralnetworks, ; fedus_switch_2022, ; jiang2024mixtralexperts, ; do2024effectivenessdiscreterepresentationssparse, ), and Expert Choice, where tokens are assigned to specific experts (zhou_mixture_experts_2022, ).
+Token Choice treats all tokens equally, which has raised concerns among researchers (wu2021attentionneed, ; hou2022tokendroppingefficientbert, ; lin2025rho1tokensneed, ), while Expert Choice suffers from token-dropping issues.
 Additionally, SMoE faces the challenge of representation collapse, where experts produce similar outputs.
-Various solutions have been proposed, such as XMoE, which employs low-dimensional routing scores (Chi et al., 2022b), and SMoE-dropout, which gradually activates more experts (Chen et al., 2023a).
-Other approaches, including HyperRouter (Do et al., 2023a) and StableMoE (Dai et al., 2022a), focus on enhancing router stability and robustness.
-Although these advancements have improved SMoE models, representation collapse remains a persistent issue (Pham et al., 2024a; Do et al., 2024a).
+Various solutions have been proposed, such as XMoE, which employs low-dimensional routing scores (chi2022representation, ), and SMoE-dropout, which gradually activates more experts (chen2023sparse, ).
+Other approaches, including HyperRouter (do2023hyperrouter, ) and StableMoE (dai2022stablemoe, ), focus on enhancing router stability and robustness.
+Although these advancements have improved SMoE models, representation collapse remains a persistent issue (pham2024competesmoe, ; do2024simsmoesolvingrepresentationalcollapse, ).
 Our approach addresses this by optimizing the alignment between tokens and the most suitable experts, expanding expert specialization and mitigating collapse.
-Competitive Learning.
-Competitive learning is an unsupervised learning approach where computational units compete to respond to a given input, enabling feature discovery (McClelland and Rumelhart, 1987; Andersen et al., 1969; Stefanis and Jasper, 1969).
-Inspired by the biological brain, this concept has recently been leveraged to enhance the efficiency of Large Language Models (LLMs) (Cai et al., 2024; Zhao et al., 2024).
-Furthermore, competitive learning has been shown to enhance the efficiency of Mixture of Experts (MoE) models (Ahn and Sentis, 2021; Cai et al., 2023).
-The most relevant work to our work is CompeteSMoE (Pham et al., 2024b), which frames competition through expert outputs and employs a router trained to predict competition outcomes in a scheduled manner.
-However, it shares a key limitation with traditional competitive learning frameworks, activating all experts, which becomes infeasible for large-scale models with millions of parameters.
-In contrast, we redefine competition among expert embeddings into two categories: Token Choice SMoE, which represents horizontal competitive learning, and Expert Choice SMoE, which represents vertical competitive learning.
-Building on this perspective, we introduce two novel scoring methods and mechanisms - Unified Competitive Score and Unified Competitive Mechanism - that enhance traditional SMoE in both performance and efficiency.
+Moreover, we also discuss related work on Linear Programming and other studies relevant to our approach in Section A.1.
 3 Methodology
-We introduce Unified Competitive Learning SMoE (USMoE), a novel and efficient Sparse Mixture of Experts framework designed to address the limitations of both Token Choice and Expert Choice through a unified competitive learning mechanism.
-As shown in Section 3.2, our approach consists of two key components:
-  • Unified Competitive Score – a scoring function that balances expert selection.
-  • Unified Competitive Mechanism – a structured routing strategy that ensures efficient and effective expert allocation.
+We introduce Unified Sparse Mixture of Experts (USMoE), a novel and efficient Sparse Mixture of Experts framework designed to address the limitations of both Token Choice and Expert Choice through a Unified Score, a scoring function that balances expert and token important and Unified Mechanism, a structured routing strategy that consider information from both token dimension and expert dimension.
 3.1 Preliminaries
-Sparse Mixture of Experts.
-The Sparse Mixture of Experts (SMoE) is typically a transformer architecture that replaces the multi-layer perceptron (MLP) layers in standard transformers with Mixture of Experts (MoE) layers, inspired by (Shazeer et al., 2017a).
-Given $\mathbf{x}\in\mathbb{R}^{b\times d}$ as the output of the multi-head attention (MHA) layer, the result of the SMoE with $n$ experts is a weighted sum of each expert’s computation, $E_{i}(x)$, weighted by the router function $\mathcal{S}(x)$:
-$f_{\mathrm{SMoE}}(\boldsymbol{x})=\sum_{i=1}^{n}\mathcal{S}(\boldsymbol{x})_{i}\cdot E_{i}(\boldsymbol{x})$ (1)
-Where $\mathcal{S}(x)$ is computed by $TopK$ function as below the Equation  2, and $W_{e}$ is a learnable experts embeddings.
-$\mathcal{S}(\boldsymbol{x})=\operatorname{TopK}(\operatorname{softmax}(W_{e}\cdot x),k)$ (2)
-$\operatorname{TopK}(\boldsymbol{v},k)=\begin{cases}\boldsymbol{v_{i}}&\text{if}\boldsymbol{v_{i}}\text{ is in the top }k\text{ largest of }\boldsymbol{v},\\ -\infty&\text{otherwise}.\end{cases}$
-3.2 Unified Competitive Learning SMoE (USMoE)
-Unified Competitive Score.
-As illustrated in Figure 2, the Token Choice approach selects and promotes the best expert by comparing the similarity scores between a token and the available experts.
-Given a representation $\mathbf{x}\in\mathbb{R}^{b\times l\times d}$, which is the output of the multi-head attention (MHA) layer, and a router matrix $W_{e}\in\mathbb{R}^{d\times n}$, where $n$ represents the number of experts, the similarity score $s_{t}$ in the Token Choice approach is computed as:
-$s_{t}=\operatorname{softmax}(W_{e}\cdot x,d=-1)$ (3)
-where $d$ denotes the dimension along which the softmax function is applied.
-In contrast, the Expert Choice approach selects and promotes the best tokens by comparing the similarity scores between an expert and the available tokens.
-The similarity score $s_{e}$ in the Expert Choice approach is computed as:
-$s_{e}=\operatorname{softmax}(W_{e}\cdot x,d=1)$ (4)
-While Expert Choice ensures an equal distribution of tokens across experts, Token Choice guarantees that each token is fairly processed by the model.
-To leverage the strengths of both approaches, we introduce the Unified Competitive Score, a soft combination of the Token Choice and Expert Choice scores, defined as:
-$s_{u}=\alpha\cdot s_{e}+(1-\alpha)\cdot s_{t}$ (5)
-where $\alpha\in[0,1]$ is a tunable control parameter that can be adjusted based on the data. In practice, we find that $\alpha\approx 0.5$ is an appropriate choice.
-Unified Competitive Mechanism.
-Token Choice selects experts using horizontal competitive learning, where each token is assigned to the most similar expert.
-This approach performs well when there is a highly relevant expert for a given token.
-However, its effectiveness diminishes when the similarity scores between the token and all experts are low.
-In contrast, Expert Choice selects tokens using vertical competitive learning, where each expert is assigned the most similar tokens.
-This approach is effective when the similarity score distribution across experts is well-structured.
-However, this condition is difficult to achieve due to the issue of representation collapse  Chi et al.
-(2022b).
-As a result, Expert Choice may either discard important tokens or lead to multiple experts selecting the same token.
-USMoE addresses these challenges by introducing the Unified Competitive Mechanism, which treats expert and token selection as a joint (expert, token) pairing process.
-As demonstrated in Algorithm  1, we first flatten the similarity matrix and then select the top-$N$ highest-scoring pairs, representing the most similar expert-token combinations.
-Let the ordered array $(S^{t}_{1},S^{t}_{2},\dots,S^{t}_{N})$ represent the top-$N$ token scores selected using the Token Choice approach, where $S^{t}_{1}\leq S^{t}_{2}\leq\dots\leq S^{t}_{N}$.
-Similarly, let $(S^{e}_{1},S^{e}_{2},\dots,S^{e}_{N})$ denote the top-$N$ token scores selected using the Expert Choice approach, where $S^{e}_{1}\leq S^{e}_{2}\leq\dots\leq S^{e}_{N}$.
-Finally, let $(S^{u}_{1},S^{u}_{2},\dots,S^{u}_{N})$ represent the top-$N$ token scores selected using the Unified Competitive Score approach, where $S^{u}_{1}\leq S^{u}_{2}\leq\dots\leq S^{u}_{n}$.
-We derive the following inequalities, with a detailed proof provided in the Appendix A.1:
-$S^{t}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N]$ (6)
-$S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N]$
-Based on Inequality 6, the Unified Competitive Mechanism ensures that the selected expert-token pairs are at least as optimal as those chosen by the Token Choice approach.
-Moreover, this mechanism addresses the limitations of Expert Choice, which may drop important tokens, and Token Choice, which may select irrelevant tokens.
-Figure 2: An illustration of our USMoE framework, which enhances existing SMoE models through a Unified Competitive Learning approach.
-The method first flattens the score matching between tokens and experts into a 1D representation, then selects the Top-N best pairs, and finally maps the results back to the original 2D matching scores.
+The Sparse Mixture of Experts (SMoE) architecture replaces the MLP layers in standard transformers with Mixture of Experts (MoE) layers (shazeer2017outrageously, ).
+Let $h\in\mathbb{R}^{T\times d}$ denote the token representations from the attention layer, where $T$ is the number of tokens and $d$ is the hidden dimension.
+Given $N$ expert functions $\{\text{FFN}_{j}\}_{j=1}^{N}$, a gating projection matrix $W\in\mathbb{R}^{d\times N}$ maps tokens to expert affinity scores.
+Denotes $S=f(hW)\in\mathbb{R}^{T\times N}$ is a compatibility score, where $f$ maps the scores to a routing distribution (e.g., softmax or sigmoid).
+Given $X\in\{0,1\}^{T\times N}$ be a binary routing matrix, where $x_{ij}=1$ if token $i$ is routed to expert $j$, the output of the SMoE layer is:
+$f_{\text{SMoE}}(h)=\sum_{j=1}^{N}S_{\cdot j}\odot X_{\cdot j}\odot\text{FFN}_{j}(h\odot X_{\cdot j}),$ (1)
+where $\text{FFN}_{j}(\cdot)$ denotes the computation performed by expert $j^{th}$, and $\odot$ represents element-wise multiplication, and FFN denotes for the Feed-forward neural networks.
+3.2 Experts Selection as Linear Programming
+Viewing expert selection as a linear programming problem allows for globally optimizing expert assignments by maximizing similarity scores under a strict sparsity (computational) constraint.
+Expert Selection Optimization.
+Given a computational budget of $c$ experts per token, the objective of Sparse Mixture-of-Experts (MoE) routing is to determine the optimal expert assignment $X^{*}$ that maximizes the total compatibility score, subject to a sparsity constraint $c$:
+maximize $\sum_{i=1}^{T}\sum_{j=1}^{N}S_{ij}x_{ij}$ (2)
+subject to $\sum_{i=1}^{T}\sum_{j=1}^{N}x_{ij}\leq c$
+$x_{ij}\in\{0,1\},\quad\forall i,j.$
+Figure 2: An illustration of our USMoE selection mechanism (middle) is shown, which incorporates information from both the token and expert dimensions.
+In contrast, Token Choice (TC, left) considers only the expert dimension, while Expert Choice (EC, right) focuses solely on the token dimension.
+Tokens marked with # represent noisy or irrelevant tokens.
+TC struggles to handle these # tokens effectively, and EC is prone to missing important tokens.
+USMoE addresses both issues, making our method more robust than traditional MoE approaches.
+This robustness is demonstrated in both theoretical analysis and experimental results.
+’E’ denotes experts.
 Best viewed in color.
+Proposition 3.1.
+Let $S\in\mathbb{R}^{T\times N}$ be the compatibility score matrix and $c\in\mathbb{N}$ a global routing budget. Consider the objective:
+$M=\langle S,X\rangle=\sum_{i=1}^{T}\sum_{j=1}^{N}S_{ij}x_{ij},$
+subject to the constraint $\sum_{i,j}x_{ij}\leq c$, with $x_{ij}\in\{0,1\}$.
+Let $X_{\text{USMoE}}=\text{TopK}(S,c)$ be the binary mask produced by selecting the top-$c$ entries of $S$ globally.
+Then for any other feasible binary routing matrix $X_{T}\in\{0,1\}^{T\times N}$ satisfying the same budget constraint $\sum_{i,j}(X_{T})_{ij}\leq c$, we have:
+$\langle S,X_{T}\rangle\leq\langle S,X_{\text{USMoE}}\rangle.$
+The Proposition 3.1 indicate that the Unified Mechanism (USMoE) yields the optimal solution that maximizes the total similarity under the global budget constraint.
+The proof is provided in Appendix  A.
+Definition 3.2 (Unified Score Function).
+Let $S\in\mathbb{R}^{T\times N}$ be the compatibility score matrix between $T$ tokens and $N$ experts. Define two score mapping functions:
+  • $f_{t}(S)\in\mathbb{R}^{T\times N}$: a row-wise scoring function used in Token Choice (e.g., softmax applied across each row).
+  • $f_{e}(S)\in\mathbb{R}^{T\times N}$: a column-wise scoring function used in Expert Choice (e.g., softmax applied across each column).
+The Unified Score Function combines both perspectives via a linear combination:
+$f_{\text{USMoE}}(S)=\alpha\cdot f_{e}(S)+\beta\cdot f_{t}(S),$
+where $\alpha,\beta\in\mathbb{R}$ are non-negative coefficients such that $\alpha+\beta=1$.
+The Unified Score as Definition 3.2 integrates both token-centric and expert-centric preferences to inform more balanced routing decisions.
+Lemma 3.3.
+The Unified Score Function $f_{\text{USMoE}}(S)=\alpha\cdot f_{e}(S)+\beta\cdot f_{t}(S)$, where $\alpha,\beta\in\mathbb{R}$ are non-negative coefficients such that $\alpha+\beta=1$, is more robust to representation collapse chi2022representation than using $f_{t}(S)$ or $f_{e}(S)$ alone.
+Lemma 3.3 establishes that USMoE outperforms conventional SMoEs by mitigating the representation collapse issue, as discussed in chi2022representation . A detailed proof is presented in Appendix A.
+3.3 Unified Sparse Mixture of Experts (USMoE)
+Revising Expert Choice.
+The Expert Choice (EC) approach faces scalability challenges in autoregressive models due to information leakage wang2024auxiliarylossfreeloadbalancingstrategy ; raposo2024mixtureofdepthsdynamicallyallocatingcompute introduced by the softmax operator.
+To address this, we replace softmax with a sigmoid function for Token Choice, which eliminates inter-token leakage and, as shown in Figure 6, leads to improved bits-per-character (BPC) performance and a reduced token dropping ratio.
+Finding 1: The Sigmoid function outperforms Softmax in Expert Choice MoE, as it handles information leakage and token dropping more effectively.
+Unified Score.
+We empirically observe that Token Choice effectively captures in-context learning behavior, while Expert Choice conveys rich semantic alignment as Figure 3.
+To harness the complementary strengths of both, we propose the Unified Score, defined in Algorithm 1, as a weighted sum of the Token Choice and Expert Choice scores.
+This unified formulation integrates both token-centric and expert-centric signals to enable more balanced and informed routing decisions.
+Empirically, we find that setting the combination weight to $\alpha\approx 0.5$ yields a robust trade-off between the two strategies.
+Figure 3: The performance of Unified Score (USMoE), Softmax (TC), and Sigmoid (EC) across the MTEB benchmark.
+Sigmoid outperforms Softmax on tasks without prompting, indicating stronger semantic representations.
+Finding 2: Sigmoid (EC) excels at capturing semantic information, while Softmax (TC) is better suited for leveraging in-context cues.
+Unified Mechanism.
+We propose the Unified Mechanism, which formulates expert-token assignment as a joint selection problem.
+By flattening the similarity matrix and selecting the top-$N$ expert-token pairs based on the highest unified scores, as illustrated in Figure 3.1, this mechanism enables efficient, context-aware routing within sparse MoE architectures.
+[…]
+2:Output $Y$
+[…]
+[…]
+[…]
+[…]
+[…]
+8:return $Y$
 Algorithm 1 USMoE Layer
-1:$X\in\mathbb{R}^{B\times L\times D}$, router weights $R\in\mathbb{R}^{D\times N}$, experts, controlling factor $\alpha$
-2:Output $Y$
-3:Compute logits:
-4:      $\text{logits}\leftarrow X\cdot R$triangleright Dot product Similarity
-5:Compute token choice score:
-[…]
-7:Compute expert choice score:
-[…]
-9:Compute unified score:
-[…]
-11:Reshape score:
-12:      $U\leftarrow\operatorname{reshape}(U,B,-1)$
-13:Compute top-$N$ indices and values:
-[…]
-15:      $\operatorname{TopK}(U,\text{topn},\text{dim}=1)$
-16:Compute output using SMoE:
-17:      $Y\leftarrow$
-[…]
-19:return $Y$
-3.3 Theoretical Analysis for Representation Collapse of Sparse Mixture of Experts
-Following  (Chi et al., 2022a) and  (Do et al., 2023b), we illustrate the representation collapse issue using the Jacobian matrix approach.
+4 Experiments
+In this section, we evaluate our method across both Large Language Models (LLMs) and Vision tasks, under clean and adversarial (attack) settings, and in both training-free and fine-tuned scenarios.
+We empirically demonstrate the advantages of USMoE over Token Choice (TC) and Expert Choice (EC) across advanced Sparse Mixture of Experts (SMoE) models, including QwenMoE qwen_moe , OLMoE muennighoff2024olmoeopenmixtureofexpertslanguage , and DeepSeekMoE dai2024deepseekmoeultimateexpertspecialization .
+Through extensive experiments, we show that: (1) USMoE outperforms baseline methods even without additional computational cost; (2) USMoE provides significant improvements across both language and vision domains; (3) Our method demonstrates robustness in both LLMs and vision tasks; and (4) USMoE supports flexible Top-k expert selection, which is valuable for scenarios with limited computational resources while maintaining competitive performance.
+4.1 Large Language Models (LLMs)
+We evaluate our method against two standard SMoE baselines across the following three settings:
+Training-Free.
+We evaluate our approach on three state-of-the-art Sparse Mixture of Experts (MoE) models: (1) OLMoE-1B-7B muennighoff2024olmoeopenmixtureofexpertslanguage , which has 7B parameters, 16 layers, and 64 experts per layer; (2) Qwen1.5-MoE-A2.7B qwen_moe , comprising 7B parameters, 24 layers, and 60 experts per layer; and DeepSeekMoE-16B which consists of 16B parameters, 28 layers, and 64 experts per layer.
+We evaluate these models on the Massive Text Embedding Benchmark (MTEB) muennighoff2023mtebmassivetextembedding without additional fine-tuning.
+Supervised Fine-Tuning.
+We demonstrate the effectiveness of USMoE on supervised fine-tuning using the Alpaca dataset alpaca , adopting the memory-efficient full-parameter training approach proposed by zhao2024galorememoryefficientllmtraining , which outperforms common low-rank adaptation methods in terms of memory efficiency.
+To evaluate the robustness of our method, we follow the procedure in nielsen2025tightclustersmakespecialized , applying a Text Attack strategy where words are randomly replaced with a generic token “AAA”.
+Training from scratch.
+To assess the effectiveness of our method, we compare USMoE with the Token Choice approaches, including SMoE jiang2024mixtralexperts , SMoE-Dropout (abbreviated as "SMoE-DR"), XMoE chi2022representation , and StableMoE dai2022stablemoe , as well as the Expert Choice approach zhou2022mixtureofexpertsexpertchoicerouting for pre-training and fine-tuning tasks.
+We follow the approach of (chen_sparse_2023, ) and use a base Transformer-XL dai_transformer_xl_2019 with four decoder layers.
+We train both base and large-scale versions of Transformer-XL on four datasets (Enwik8, Text8, Wikitext-103, and One Billion Words) for 100k iterations, following the implementation in chen_sparse_2023 .
+Then we fine-tune the pre-trained weights for text classification tasks, including SST-2 socher_recursive_2013 , SST-5 socher_recursive_2013 , IMDB maas_learning_2011 , and BANKING77 casanueva-etal-2020-efficient .
+More implementation details and additional results are provided in the Appendix A.
+4.1.1 Training-Free
+In this section, inspired by  li2024mixtureofexpertsllmsecretlyembedding , we test our method as a plug-in framework on well-trained SMoE models, including OLMoE-1B-7B muennighoff2024olmoeopenmixtureofexpertslanguage :, DeepSeekMoE-16B  dai2024deepseekmoeultimateexpertspecialization , Qwen1.5-MoE-A2.7B qwen_moe .
+We evaluate performance on a subset of tasks from the Massive Text Embedding Benchmark (MTEB)  muennighoff2023mtebmassivetextembedding , which covers key downstream applications for sentence embeddings, including Classification, Clustering, Pair Classification, Re-ranking, Retrieval, Semantic Textual Similarity (STS), and Summarization.
+Following the MTEB evaluation framework, we use Accuracy for Classification, V-Measure for Clustering, Average Precision for Pair Classification, Mean Average Precision for Re-ranking, nDCG for Retrieval, and Spearman’s correlation for STS and Summarization.
+Our method consistently demonstrates performance improvements across a range of MTEB tasks .
+Detailed results for datasets under each task type are provided in Appendix A.
+USMoE outperforms both the Expert Choice and Token Choice approaches in most cases, underscoring the complementary nature of these two methods.
+For tasks evaluated in Figure  4, USMoE proves even more effective at enhancing the Token Choice approach, delivering notable gains of 14%, 12%, and 13% for OLMoE-1B-7B, Qwen1.5-MoE-A2.7B, and DeepSeekMoE-16B, respectively, across MTEB tasks without additional training.
+Specifically, Qwen1.5-MoE-A2.7B achieves a remarkable improvement from 13.4% (Token Choice) to 40.0% (USMoE) in the Summarization task, representing a 198% gain.
+This trend persists across DeepSeekMoE-16B and OLMoE-1B-7B, where USMoE consistently outperforms both the Token Choice and Expert Choice approaches.
+Overall, our approach outperforms the baselines in terms of performance while exhibiting lower variance across multiple tasks and different runs.
+Figure 4: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks and advance SMoE models.
+The best result for each row is highlighted in bold.
+Best viewed in color.
+4.1.2 Supervised Fine-Tuning
+Supervised Fine-Tuning.
+We demonstrate the efficiency and robustness of USMoE in supervised fine-tuning across two advanced SMoE models: QwenMoE (7B)qwen_moe and OLMoE (7B)muennighoff2024olmoeopenmixtureofexpertslanguage .
+For training, we adopt Supervised Fine-Tuning (SFT) using GaLorezhao2024galorememoryefficientllmtraining , which enables full-parameter optimization with improved memory efficiency compared to typical low-rank adaptation methods such as LoRA.
+The fine-tuning is conducted on the Alpaca datasetalpaca for 2,000 steps, with evaluation performed every 10 steps.
+Furthermore, we assess our method on both clean (referred to as QwenMoE and OLMoE) and adversarially corrupted versions (denoted QwenMoE-Corrupt and OLMoE-Corrupt), where sequences of random “AAA” tokens are injected to simulate noise.
+Robustness.
+Figure 5 presents the training results for both QwenMoE and OLMoE models.
+The results show that USMoE consistently outperforms the baselines across both QwenMoE and OLMoE models.
+Our method surpasses both Token Choice (TC) and Expert Choice (EC), particularly on QwenMoE, where it achieves significantly lower training and validation perplexity, indicating faster convergence and better generalization.
+Notably, the performance gap becomes even more pronounced under the corrupted setting, highlighting the superior robustness of USMoE compared to traditional SMoE approaches.
+Figure 5: Illustration of comparing the performance of USMoE, Token Choice (TC), Expert Choice (EC) using QwenMoE and OLMoE for the Supervised Fine-Tuning task on Apaca dataset for 2K steps under both clean and corrupted settings.
+Training and validation perplexity over training steps are reported, and lower values are better.
+4.1.3 Training from Scratch
+Robustness.
+Table 2 presents pre-training results on four datasets (enwik8, text8, WikiText-103, and One Billion Words) under both clean and corrupted settings.
+USMoE consistently outperforms both Expert Choice and Token Choice routing methods across all datasets.
+Similar to the SFT task, the performance gap widens under corruption, especially on One Billion Words—highlighting the robustness advantage of USMoE over traditional SMoEs.
+For additional experiments on training from scratch, such as scaling models up to 400M parameters or fine-tuning on downstream tasks—please refer to Section A.
+Flexibility.
+One of USMoE’s key strengths is its flexibility in computational resource usage.
+Unlike traditional SMoEs, which require the number of selected experts (top-$k$) to be an integer, USMoE supports fractional expert selection.
+This allows the model to maintain competitive performance while leveraging fewer experts.
+For instance, on text8, USMoE outperforms SMoE using only 1.5 experts.
+It also surpasses Token Choice, which uses two experts, on enwik8, text8, and One Billion Words, while averaging just 1.5 experts per token.
+Moreover, using only 1.5 experts enables USMoE to reduce FLOPs by 14% compared to SMoE, with minimal performance trade-off.
+This makes USMoE particularly well-suited for resource-constrained environments such as IoT and edge devices.
+Setting Original Corrupt
+Transformer-XL(20M) Topkk Enwik8 Text8 WikiText-103 lm1b Enwik8 Text8 WikiText-103 lm1b
+USMoE 2 1.18 1.20 29.20 56.90 1.75 1.83 38.45 68.43
+1.5 1.19 1.28 30.67 57.55 1.78 1.95 40.39 70.47
+TC 2 1.20 1.29 30.16 58.00 1.77 1.86 39.31 79.73
+EC 2 1.18 1.24 29.83 58.60 1.76 1.89 39.28 71.75
+Table 2: Performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) across multiple datasets, with BPC on the Enwik8 and Text8 test sets, and perplexity on the WikiText-103 and One Billion Word test sets.
+Lower values are better, with the best results highlighted in bold.
+4.2 Vision
+Effiency.
+To ensure a thorough and meaningful performance comparison, we evaluate against two categories of baselines that are closely aligned with USMoE, including Token Choice, Expert Choice, and SoftMoE (puigcerver2024sparsesoftmixturesexperts, ), currently among the most advanced MoE models in the vision domain.
+We explore two Mixture of Experts configurations: (1) a small model with 10 million parameters (10M), and (2) a large model with 110 million parameters (110M).
+As shown in Table 3, USMoE consistently outperforms Vision Transformer variants with Token Choice, Expert Choice, and SoftMoE across all eight tasks and four image classification datasets.
+Our experiments are conducted three times on four datasets (CIFAR-10, CIFAR-100, STL-10, SVHN, and ImageNet1K), using different random seeds.
+We report average results along with standard deviations.
+USMoE not only achieves superior average performance but also demonstrates greater stability, as evidenced by lower standard deviation values.
+ViT-MoE 10M 110M Avg.
+Dataset Cifar10 Cifar100 STL-10 SVHN ImageNet-1K Cifar10 Cifar100 STL-10 SVHN ImageNet-1K
+USMoE 89.6±0.3 66.6±0.5 66.7±0.4 95.6±0.1 60.2±0.1 91.5±0.5 67.3±0.5 66.2±0.1 96.1±0.1 73.5±0.4 77.3±0.3
+TC 88.7±0.2 65.4±0.5 66.4±0.1 95.1±0.1 56.6±0.5 85.7±0.8 55.5±2.8 64.4±0.2 94.5±0.4 72.0±0.4 74.4±1.5
+EC 88.9±0.3 65.7±0.4 66.1±0.4 95.0±0.1 56.2±0.4 87.4±0.7 66.2±0.9 65.5±0.4 93.2±0.2 70.9±0.5 75.5±1.2
+SoftMoE 85.6±0.3 61.4±0.3 65.4±0.2 94.8±0.1 46.8±0.6 80.3±1.0 42.9±1.4 63.9±1.2 93.5±0.1 71.2±0.3 70.6±1.5
+Table 3: Accuracy of models evaluated on vision datasets. Each method is evaluated 3 times, reporting the mean and standard deviation. Higher is better, the best results are in bold.
+Robustness.
+Table 4 presents a comparative analysis of ViT-MoE variants, USMoE, Token Choice (TC), Expert Choice (EC), and SoftMoE, under three common adversarial attacks: Projected Gradient Descent(PGD)(chang2018efficienttwostepadversarialdefense, ), Fast Gradient Sign Method(FGSM) goodfellow2015explainingharnessingadversarialexamples , and Simultaneous Perturbation Stochastic Approximation (SPSA) (SPALL1997109, ), across five standard image classification datasets.
+USMoE consistently achieves strong performance, outperforming or matching the best baselines in most scenarios.
+Notably, it achieves the highest average robustness across all three attack types, particularly excelling under PGD (46.2%), FGSM (43.3%), and SPSA (63.9%).
+While Token Choice (TC) performs poorly across all three attack types, USMoE consistently delivers stronger and more reliable performance on average.
+These results highlight USMoE’s robustness and generalization capability under adversarial conditions, making it a more stable and effective choice for secure image classification tasks.
+Finding 3: Token Choice (TC) struggles under adversarial attacks.
+Dataset PGD FGSM SPSA
+ViT-MoE (10M) USMoE TC EC SoftMoE USMoE TC EC SoftMoE USMoE TC EC SoftMoE
+CIFAR‑10 57.4 27.7 57.5 55.6 49.5 26.4 48.5 45.8 83.6 33.4 87.0 69.1
+CIFAR‑100 27.5 11.9 27.0 28.6 21.5 11.1 19.3 20.6 65.8 15.8 59.1 39.3
+STL‑10 41.0 37.7 39.5 39.8 36.5 35.8 34.8 35.0 64.3 44.1 61.3 49.6
+SVHN 81.0 38.4 80.5 91.3 76.0 36.1 75.4 69.2 92.5 36.3 91.8 84.4
+ImageNet‑1K 24.0 1.3 22.4 13.3 12.2 1.4 11.2 9.9 13.4 1.5 12.3 10.9
+Avg. 46.2 23.4 45.4 45.7 43.3 22.2 37.8 36.1 63.9 26.2 62.3 50.7
+Table 4: Robustness evaluation of different ViT-MoE models under adversarial attacks: PGD, FGSM, and SPSA across five datasets. Bold indicates the best performance for each task and attack.
+5 Conclusion
+In this work, we reinterpret Token Choice and Expert Choice Sparse Mixture of Experts (SMoE) through a linear programming lens, revealing their inherent limitations.
+Based on this analysis, we propose Unified Sparse Mixture of Experts (USMoE) - a novel and efficient framework that advances SMoE by introducing a unified mechanism and unified score learning.
+We theoretically prove that USMoE achieves superior expert selection compared to traditional methods, effectively improving expert learning capacity while mitigating expert collapse.
+As a result, USMoE learns more robust expert representations and overcomes the representation collapse issues commonly observed in conventional SMoE training.
+Extensive experiments across diverse domains, including vision and large language models (LLMs), under both clean and adversarial settings (covering training-free and fine-tuning scenarios), demonstrate that USMoE enables more efficient and effective training and inference compared to state-of-the-art routing strategies.
+Appendix A Appendix
+Supplementary Material for “Unified Sparse Mixture of Experts”
+This document is organized as follows: Appendix A.1 discuss related work on Linear Programming and other studies relevant to our approach.
+We also discuss more about the Ablation Studies at Appendix A.2.
+Appendix A.3 provides the theoretical proof supporting the discussion in Section 3.
+Following this, Appendix A.4 additional experiment results that show the advantage of our method.
+Appendix A.5 offers a detailed analysis of the Token Choice and Expert Choice routers and explains why our method outperforms the baseline approaches.
+Appendix A.6 describes the implementation details in full.
+A.1 Related Work
+Linear Programming.
+Linear programming, a foundational method in mathematical optimization, has gained relevance in the domain of Large Language Models (LLMs), where task allocation and dependency management among skills can be effectively modeled as constrained optimization problems [40, 33].
+Recent work such as [22] introduces an Integer Linear Programming (ILP) formulation to jointly optimize expert placement by accounting for token load, communication overhead, and computation cost.
+Framing SMoEs from the perspective of linear programming not only offers deeper theoretical understanding but also opens promising directions for developing more efficient and principled routing strategies.
+The most relevant work to ours is Expert Race[56], which proposes a flexible routing strategy for diffusion transformers.
+However, similar to the traditional Expert Choice (EC) approach, it suffers from a critical limitation, information leakage - where a model unintentionally accesses future tokens [43, 54], which becomes especially problematic in autoregressive language modeling settings[43, 54].
+In contrast, we propose a modified version of Expert Choice [58] by replacing the softmax mapping function with a sigmoid function and limiting comparisons to within the token dimension.
+This design effectively mitigates information leakage, making Expert Choice more suitable for large language models.
+Additionally, we introduce two novel modules: (1) the Unified Score, a linear combination of the mapping functions from Token Choice and Expert Choice, and (2) the Unified Mechanism, which jointly considers both expert and token dimensions when selecting experts.
+Both theoretical analysis and empirical results demonstrate that our method outperforms traditional SMoE approaches in terms of both performance and robustness.
+A.2 Ablation Studies
+We investigate the effectiveness and robustness of USMoE to the different hyper-parameter settings.
+A.2.1 Unified Learning Strategy Comparison
+The Unified Mechanism is implemented using two approaches: (1) a sequence-based method that compares all tokens within a sequence (referred to as "USMoE-Sequence") and (2) a batch-based method that compares all tokens within a batch or mini-batch (referred to as "USMoE-Batch").
+We evaluate both approaches on the Classification task, with results presented in Table 10.
+The findings indicate that both methods outperform the Expert Choice and Token Choice approaches, demonstrating the effectiveness of our method.
+Notably, the sequence-based approach achieves superior performance in the Classification task, as it ensures that no important tokens are missed within a sequence - an assurance that the batch/mini-batch implementation may not always provide.
+A.2.2 Robustness to the controlling factor
+The Unified Score ($\alpha$) enables the model to adjust its scoring mechanism, either favoring a diverse set of experts per sequence or distributing experts more evenly across tokens.
+We evaluate the robustness of the controlling factor $\alpha$ on the classification task using the DeepSeekMoE-16B model, with results presented in Table 11.
+When $\alpha=0.0$, the scoring mechanism aligns with Token Choice, while at $\alpha=1.0$, it follows Expert Choice.
+Overall, USMoE demonstrates strong effectiveness within the range of $\alpha\in(0.3,0.7)$, striking a balance between expert diversity and token importance.
+This range provides an optimal trade-off between enforcing SMoE’s unified policy and enhancing traditional approaches for the task, leading to superior overall performance.
+Notably, all tested $\alpha$ configurations outperform the Expert Choice approach ($\alpha=1.0$).
+A.3 Theoretical Proof for Section  3
+Definition A.1 (General Routing Function $\mathcal{R}$).
+Let $S\in\mathbb{R}^{T\times N}$ be the compatibility score matrix between $T$ tokens and $N$ experts, and let $c\in\mathbb{N}$ denote a global routing budget that constrains the total number of token-to-expert assignments.
+Let $\text{TopK}_{R}$, $\text{TopK}_{C}$, and TopK denote the TopK selection operators applied row-wise, column-wise, and globally, respectively.
+We define a general routing function $\mathcal{R}(S,c,\texttt{mode})$ that constructs a sparse binary routing matrix $X\in\{0,1\}^{T\times N}$, where $x_{ij}=1$ indicates that token $i$ is routed to expert $j$, as follows:
+$X=\mathcal{R}(S,c,\texttt{mode})=\begin{cases}\text{TopK}(S,c),&\texttt{mode}=\text{USMoE},\\[6.0pt] \text{TopK}_{R}(S,\left\lfloor\frac{c}{T}\right\rfloor),&\texttt{mode}=\text{Token Choice},\\[6.0pt] \text{TopK}_{C}(S,\left\lfloor\frac{c}{N}\right\rfloor),&\texttt{mode}=\text{Expert Choice},\end{cases}$
+Lemma A.2.
+The Unified Mechanism (USMoE), as defined in Definition A.1, yields a superior solution to the optimization problem in Equation 2 compared to both Token Choice and Expert Choice mechanisms.
+The proof is given in Appendix A, where Lemma A.2 establishes the superiority of USMoE over Token Choice and Expert Choice.
+A.3.1 Proof of Lemma  A.2
+Let the ordered arrays $(S^{t}_{1},S^{t}_{2},\dots,S^{t}_{T})$, $(S^{e}_{1},S^{e}_{2},\dots,S^{e}_{T})$, and $(S^{u}_{1},S^{u}_{2},\dots,S^{u}_{T})$ represent the top-$T$ compatibility scores selected by the Token Choice, Expert Choice, and Unified Score mechanisms, respectively.
+These scores are sorted in ascending order such that
+$S^{t}_{1}\leq S^{t}_{2}\leq\dots\leq S^{t}_{T},\quad S^{e}_{1}\leq S^{e}_{2}\leq\dots\leq S^{e}_{T},\quad S^{u}_{1}\leq S^{u}_{2}\leq\dots\leq S^{u}_{T}.$
+We claim the following inequality holds:
+$S^{t}_{i}$ $\leq S^{u}_{i},\quad\forall i\in[1,T],$ (3)
+$S^{e}_{i}$ $\leq S^{u}_{i},\quad\forall i\in[1,T].$
+Proof.
+Let $x\in\mathbb{R}^{T\times d}$ denote a matrix of token embeddings and $\mathcal{E}\in\mathbb{R}^{d\times N}$ the matrix of expert embeddings, where $d$ is the embedding dimension, $T$ is the number of tokens, and $N$ is the number of experts.
+The similarity matrix $S\in\mathbb{R}^{T\times N}$ between tokens and experts is computed via a dot product:
+$S=x\cdot\mathcal{E}.$ (4)
+Since the softmax function is monotonic, the top-$k$ selection remains invariant under softmax transformation:
+$\operatorname{TopK}(\operatorname{softmax}(S),k)=\operatorname{TopK}(S,k).$ (5)
+For the Token Choice approach, the top similarity score for each token is defined as:
+$S^{t}_{j}=\max_{k\in[1,N]}S_{jk},$ (6)
+and across all tokens, the top-$T$ scores are aggregated into $\{S^{t}_{i}\}_{i=1}^{T}$.
+In contrast, the Unified Mechanism selects the top-$T$ scores globally across all token-expert pairs:
+$S^{u}_{j}=\max_{(q,k)\in[1,T]\times[1,N]}S_{qk}.$ (7)
+By construction, it follows that:
+$S^{t}_{i}\leq S^{u}_{i},\quad\forall i\in[1,T],$ (8)
+since Token Choice selects top experts per token independently, while Unified Score considers all token-expert pairs jointly.
+Similarly, in the Expert Choice approach, each expert selects its most compatible token:
+$S^{e}_{i}=\max_{q\in[1,T]}S_{qi},\quad\forall i\in[1,N],$ (9)
+and again, the top-$T$ scores across experts are extracted as $\{S^{e}_{i}\}_{i=1}^{T}$. As the Unified approach considers the global maximums, we have:
+$S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,T].$ (10)
+Combining (8) and (10), we conclude:
+$S^{t}_{i}\leq S^{u}_{i},\quad S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,T],$ (11)
+which completes the proof. ∎∎
+A.3.2 Proof of Proposition 3.1
+Proof.
+Following  [9] and  [18], we illustrate the representation collapse issue using the Jacobian matrix approach.
 Specifically, the Jacobian matrix of the SMoE with respect to $x\in\mathbb{R}^{b\times d}$ is given as:
@@
 where $\boldsymbol{c}_{j}=\mathcal{S}(x)_{k}\left(\delta_{kj}-S_{j}\right)\boldsymbol{E}(x)_{i}$.
-The first part of Equation  7, $\mathcal{S}(x)_{k}\boldsymbol{J}^{\mathrm{FFN}}$, represents the contribution from the input token and experts to the final output.
+The first part of Equation  12, $\mathcal{S}(x)_{k}\boldsymbol{J}^{\mathrm{FFN}}$, represents the contribution from the input token and experts to the final output.
 The second part, (2) $\sum_{j=1}^{n}\boldsymbol{c}_{j}\boldsymbol{e}_{j}^{\top}$ relates to learning an improved gating function to minimize task loss.
-Furthermore, Equation  7 should be updated as a linear combination of expert embeddings.
+Furthermore, Equation  12 should be updated as a linear combination of expert embeddings.
 Due to $n<<d$ in practice, the above equation illustrates the representation.
 Given $\mathcal{S}(x)=\alpha\times\mathcal{S}_{e}(x)+(1-\alpha)\times\mathcal{S}_{t}(x)$, where $\mathcal{S}_{e}(x)$ and $\mathcal{S}_{t}(x)$ represent the similarity score functions for expert choice and token choice, respectively, the Jacobian matrix of USMoE with respect to $x\in\mathbb{R}^{b\times d}$ is expressed as:
-[…]
+$\boldsymbol{J_{U}}$ $=\mathcal{S}(x)_{k}\boldsymbol{J}^{\mathrm{FFN}}+\sum_{j=1}^{n}\mathcal{S}_{e}(x)_{k}\left(\delta_{kj}-\mathcal{S}_{e}(x)_{j}\right)\boldsymbol{E}(x)_{k}\boldsymbol{e}_{j}^{\top}$
+$\quad+\sum_{j=1}^{n}\mathcal{S}_{t}(x)_{k}\left(\delta_{kj}-\mathcal{S}_{t}(x)_{j}\right)\boldsymbol{E}(x)_{k}\boldsymbol{e}_{j}^{\top}.$
 $\Longrightarrow\boldsymbol{J_{U}}=J_{1}+\sum_{j=1}^{n}c_{j}e_{j}^{\top}+\sum_{j=1}^{n}d_{j}e_{j}^{\top}$ (13)
@@
 $d_{j}=\mathcal{S}_{t}(x)_{k}\left(\delta_{kj}-\mathcal{S}_{t}(x)_{j}\right)\boldsymbol{E}(x)_{k}.$ (16)
-Similar to the Jacobian matrix of SMoE as Equation  7, the Jacobian matrix of USMoE also consists of two terms: (1) $J_{1}$, which depends on the input token and experts for the final output; and (2) $\sum_{j=1}^{2n}o_{j}\boldsymbol{e}_{j}^{\top}$ indicates to learn better gating function to minimize the task loss.
+Similar to the Jacobian matrix of SMoE as Equation  12, the Jacobian matrix of USMoE also consists of two terms: (1) $J_{1}$, which depends on the input token and experts for the final output; and (2) $\sum_{j=1}^{2n}o_{j}\boldsymbol{e}_{j}^{\top}$ indicates to learn better gating function to minimize the task loss.
 Since $2n>>n$, USMoE is more effective than SMoE in addressing the representation collapse issue.
-4 Experiments
-We evaluate our method in two scenarios: (1) without additional training and (2) with training.
-For the first scenario, inspired by  Li and Zhou (2024), we test our method as a plug-in framework on well-trained SMoE models, including OLMoE-1B-7B Muennighoff et al.
-(2024):, DeepSeekMoE-16B  Dai et al.
-(2024), Qwen1.5-MoE-A2.7B Team (2024).
-We evaluate performance on a subset of tasks from the Massive Text Embedding Benchmark (MTEB)  Muennighoff et al.
-(2023), which covers key downstream applications for sentence embeddings, including Classification, Clustering, Pair Classification, Re-ranking, Retrieval, Semantic Textual Similarity (STS), and Summarization.
+A.3.3 Proof of Lemma 3.3
+Proof.
+Let $S\in\mathbb{R}^{t\times n}$ be the similarity matrix, and let $c\in\mathbb{N}$ be a global routing budget. The goal is to find a binary matrix $X\in\{0,1\}^{t\times n}$ such that
+$\sum_{i=1}^{t}\sum_{j=1}^{n}x_{ij}\leq c$
+and the total score
+$M=\langle S,X\rangle=\sum_{i=1}^{t}\sum_{j=1}^{n}S_{ij}x_{ij}$
+is maximized.
+Let $X_{\text{USMoE}}=\text{TopK}(S,c)$ denote the binary matrix that marks the positions of the top-$c$ largest entries in $S$.
+By construction, $X_{\text{USMoE}}$ sets exactly $c$ entries of $S$ with the highest values to 1 and all others to 0.
+Suppose there exists another feasible solution $X_{T}\in\{0,1\}^{t\times n}$ with
+$\sum_{i,j}(X_{T})_{ij}\leq c$
+and assume for contradiction that
+$\langle S,X_{T}\rangle>\langle S,X_{\text{USMoE}}\rangle.$
+Since $X_{T}$ selects at most $c$ entries from $S$, and the values of the selected entries are summed to compute $\langle S,X_{T}\rangle$, the only way for $\langle S,X_{T}\rangle$ to exceed $\langle S,X_{\text{USMoE}}\rangle$ is if some entries selected by $X_{T}$ are larger than those selected by $X_{\text{USMoE}}$.
+However, this contradicts the definition of $X_{\text{USMoE}}$ as containing the top-$c$ largest values in $S$.
+Therefore, for any feasible $X_{T}$, we must have
+$\langle S,X_{T}\rangle\leq\langle S,X_{\text{USMoE}}\rangle,$
+which proves that $X_{\text{USMoE}}$ is the optimal solution. ∎
+A.4 Experimental details and additional results
+A.4.1 Which Mapping Function for Expert Choice?
+The Expert Choice (EC) approach faces scalability challenges for large language models due to the issues of information leakage [54, 43] and token dropping [42].
+We revisit this problem and identify that these issues stem from the use of the softmax operator, which introduces information leakage across tokens, and the top-k selection applied along the batch dimension.
+To address this, we propose replacing the softmax mapping function with a sigmoid function to eliminate inter-token information leakage.
+Additionally, we mitigate batch-wise leakage by applying the top-k selection only along the sequence dimension.
+Figure 6 compares the performance of EC using softmax versus sigmoid.
+The results show that the sigmoid-based EC not only achieves better bits-per-character (BPC) performance but also reduces the token dropping ratio.
+This improvement suggests that using the sigmoid mapping function has strong potential for scaling up the EC approach in large language models.
+Figure 6: Performance comparison of the Expert Choice approach using two mapping functions: sigmoid and softmax, evaluated on training performance and token dropping ratio (lower is better).
+A.4.2 Training-free additional results
+In this section, we compare methods both with and without prompts, including PromptEOL [29].
+Inspired by  [34], we also compare our method with an approach that uses the similarity score between router and expert embeddings as the hidden representation, which we refer to as "Router Embedding" or simply "Router" Additionally, we evaluate against MoEE  [34], which leverages both Router Embedding and the hidden representation of the SMoE model as embeddings.
+Model Task Router TC EC MoEE USMoE
+OLMoE-1B-7B Classification 43.1 57.7 56.2 51.7 61.4
+Clustering 16.2 24.8 26.9 23.2 32.1
+PairClassification 53.5 62.0 58.9 66.0 68.9
+Reranking 41.7 51.3 51.0 53.2 55.1
+STS 49.4 63.5 44.2 67.8 71.1
+Summarization 25.6 28.9 29.7 30.4 30.5
+Average 38.3 48.0 44.5 48.7 53.2
+Qwen1.5-MoE-A2.7B Classification 48.8 58.0 35.2 54.0 59.7
+Clustering 14.3 34.2 29.2 30.1 37.5
+PairClassification 51.9 60.5 56.0 60.3 66.6
+Reranking 41.0 46.6 45.0 51.1 56.8
+STS 48.3 50.1 50.0 64.3 69.0
+Summarization 27.0 23.0 21.9 27.3 31.0
+Average 38.6 45.4 39.6 47.9 53.4
+DeepSeekMoE-16B Classification 48.6 56.4 55.4 53.0 60.4
+Clustering 17.8 29.0 20.3 28.5 32.8
+PairClassification 57.4 59.8 53.8 63.3 67.9
+Reranking 43.8 45.7 40.9 50.6 52.4
+STS 52.8 49.0 37.1 63.4 68.1
+Summarization 29.1 24.4 25.7 29.2 30.7
+Average 41.6 44.0 38.9 48.0 52.0
+Table 5: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across across MTEB Tasks with PromptEOL [28]. The best result for each row is highlighted in bold.
+Model Task Router TC EC MoEE USMoE
+OLMoE-1B-7B Classification 41.2 43.4 44.9 41.8 51.3
+Clustering 13.7 14.7 12.0 14.5 21.0
+PairClassification 45.3 39.1 35.5 45.7 61.7
+Reranking 37.5 37.4 35.3 39.5 46.5
+STS 39.9 24.1 18.2 39.9 59.6
+Summarization 28.4 20.9 21.1 29.8 31.8
+Average 34.3 29.9 27.8 35.2 45.3
+Qwen1.5-MoE-A2.7B Classification 43.8 50.3 25.5 47.7 52.2
+Clustering 13.6 27.4 23.2 25.2 29.9
+PairClassification 45.9 46.9 43.4 51.5 56.2
+Reranking 39.6 45.3 41.6 48.5 53.1
+STS 38.8 38.0 35.6 51.8 60.7
+Summarization 28.3 13.4 15.1 31.2 40.0
+Average 35.0 36.9 30.7 42.6 48.7
+DeepSeekMoE-16B Classification 43.4 46.6 44.7 44.4 49.4
+Clustering 13.4 18.1 13.5 17.8 21.0
+PairClassification 45.5 40.9 37.1 46.1 53.5
+Reranking 38.5 38.9 35.1 42.2 45.7
+STS 37.7 26.3 23.3 40.2 51.2
+Summarization 24.9 22.0 18.5 24.4 29.9
+Average 33.9 32.1 28.7 35.9 41.8
+Table 6: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks without prompts and models. The best result for each row is highlighted in bold.
+In this section, inspired by  [34], we test our method as a plug-in framework on well-trained SMoE models, including OLMoE-1B-7B [37]:, DeepSeekMoE-16B  [12], Qwen1.5-MoE-A2.7B [52].
+We evaluate performance on a subset of tasks from the Massive Text Embedding Benchmark (MTEB)  [38], which covers key downstream applications for sentence embeddings, including Classification, Clustering, Pair Classification, Re-ranking, Retrieval, Semantic Textual Similarity (STS), and Summarization.
 Following the MTEB evaluation framework, we use Accuracy for Classification, V-Measure for Clustering, Average Precision for Pair Classification, Mean Average Precision for Re-ranking, nDCG for Retrieval, and Spearman’s correlation for STS and Summarization.
-For the second scenario, we perform language model pre-training on diverse datasets, including Enwik8, Text8 Mahoney (2011), Wikitext-103 Merity et al.
-(2017), and One Billion Words Chelba et al.
-(2014).
-To assess performance, we fine-tune the pre-trained models on various downstream benchmarks.
-4.1 Experiment Setting
-We evaluate our method against two conventional SMoE approaches: (1) the Token Choice approach and (2) the Expert Choice approach, under three different settings: (1) without additional training, (2) pre-training, and (3) fine-tuning.
-Without Training.
-We evaluate our approach on three state-of-the-art Sparse Mixture of Experts (MoE) models: (1) OLMoE-1B-7B Muennighoff et al.
-(2024), which has 7B parameters, 16 layers, and 64 experts per layer; (2) Qwen1.5-MoE-A2.7B Team (2024), comprising 7B parameters, 24 layers, and 60 experts per layer; and DeepSeekMoE-16B which consists of 16B parameters, 28 layers, and 64 experts per layer.
-We evaluate these models on the Massive Text Embedding Benchmark (MTEB) without additional fine-tuning.
-Additionally, we compare methods both with and without prompts, including PromptEOL Jiang et al.
-(2024c).
-Inspired by  Li and Zhou (2024), we also compare our method with an approach that uses the similarity score between router and expert embeddings as the hidden representation, which we refer to as "Router Embedding" or simply "Router" Additionally, we evaluate against MoEE  Li and Zhou (2024), which leverages both Router Embedding and the hidden representation of the SMoE model as embeddings.
-Pre-training.
-To assess the effectiveness of our method, we compare USMoE with the Token Choice approaches, including SMoE Jiang et al.
-(2024b), SMoE-Dropout (abbreviated as "SMoE-DR"), XMoE Chi et al.
-(2022b), and StableMoE Dai et al.
-(2022a), as well as the Expert Choice approach Zhou et al.
-(2022a) for pre-training and fine-tuning tasks.
-We follow the approach of Chen et al.
-(2023b) and use a base Transformer-XL Dai et al.
-(2019a) with four decoder layers.
-We train both base and large-scale versions of Transformer-XL on four datasets (Enwik8, Text8, Wikitext-103, and One Billion Words) for 100k iterations, following the implementation in Chen et al.
-(2023b).
-Fine-tuning.
-We fine-tune the pre-trained weights for text classification tasks, including SST-2 Socher et al.
-(2013), SST-5 Socher et al.
-(2013), IMDB Maas et al.
-(2011), and BANKING77 Casanueva et al.
-(2020).
-More implementation details and additional results are provided in the Appendix A.
-4.2 Without Training
-Our method consistently demonstrates performance improvements across a range of MTEB tasks in two scenarios: (1) with PromptEOL Jiang et al.
-(2023) (denoted as "with prompts" for brevity), as shown in Table 1, and (2) without prompts, as shown in Table 2.
+Our method consistently demonstrates performance improvements across a range of MTEB tasks in two scenarios: (1) with PromptEOL [28] (denoted as "with prompts" for brevity), as shown in Table 5, and (2) without prompts, as shown in Table 6.
 Detailed results for datasets under each task type are provided in Appendix A.
 USMoE outperforms both the Expert Choice and Token Choice approaches in most cases, underscoring the complementary nature of these two methods.
-Figure 1 illustrates the effectiveness of our method across various MTEB tasks using the OLMoE-1B-7B model, in comparison to both the Expert Choice and Token Choice approaches.
-For tasks evaluated with prompts, USMoE achieves the highest average performance across models, showing notable improvements of 4.2%, 9.1%, and 6.6% for OLMoE-1B-7B, Qwen1.5-MoE-A2.7B, and DeepSeekMoE-16B, respectively, without requiring any additional training, as detailed in Table 1.
-Notably, DeepSeekMoE-16B demonstrates a significant improvement from 50.1% (Token Choice) to 64.3% (USMoE), reflecting a 14.2% gain in the STS task.
-For tasks evaluated without prompts, USMoE proves even more effective at enhancing the Token Choice approach, delivering notable gains of 9.3%, 7.0%, and 8.0% for OLMoE-1B-7B, Qwen1.5-MoE-A2.7B, and DeepSeekMoE-16B, respectively, across MTEB tasks without additional training.
-Specifically, OLMoE-1B-7B achieves a remarkable improvement from 24.1% (Token Choice) to 47.5% (USMoE), representing a 23.4% gain.
-This trend persists across Qwen1.5-MoE and OLMoE, where USMoE consistently outperforms both the Token Choice and Expert Choice approaches.
 Interestingly, Router Embedding is less affected by prompting on the Classification dataset, as shown in Figure 8(a), while Token Choice, Expert Choice, and USMoE (ours) achieve significant performance improvements in the prompting setting.
@@
 Overall, our approach outperforms the baselines in terms of performance while exhibiting lower variance across multiple tasks and different runs.
-Model Task Router TC EC MoEE USMoE
-OLMoE-1B-7B Classification 43.1 57.7 56.2 51.7 59.2
-Clustering 16.2 24.8 26.9 23.2 30.5
-PairClassification 53.5 62.0 58.9 66.0 66.8
-Reranking 41.7 51.3 51.0 53.2 54.7
-STS 49.4 63.5 44.2 67.8 71.1
-Summarization 25.6 28.9 29.7 30.4 30.9
-Average 38.3 48.0 44.5 48.7 52.2
-Qwen1.5-MoE-A2.7B Classification 48.8 58.0 35.2 54.0 59.2
-Clustering 14.3 34.2 29.2 30.1 35.8
-PairClassification 51.9 60.5 56.0 60.3 63.4
-Reranking 41.0 46.6 45.0 51.1 53.7
-STS 48.3 50.1 50.0 64.3 66.3
-Summarization 27.0 23.0 21.9 27.3 48.4
-Average 38.6 45.4 39.6 47.9 54.5
-DeepSeekMoE-16B Classification 48.6 56.4 55.4 53.0 57.3
-Clustering 17.8 29.0 20.3 28.5 31.9
-PairClassification 57.4 59.8 53.8 63.3 65.3
-Reranking 43.8 45.7 40.9 50.6 52.1
-STS 52.8 49.0 37.1 63.4 66.0
-Summarization 29.1 24.4 25.7 29.2 30.9
-Average 41.6 44.0 38.9 48.0 50.6
-Table 1: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across across MTEB Tasks with PromptEOL Jiang et al.
-(2023).
-The best result for each row is highlighted in bold.
-Model Task Router TC EC MoEE USMoE
-OLMoE-1B-7B Classification 41.2 43.4 44.9 41.8 47.0
-Clustering 13.7 14.7 12.0 14.5 16.9
-PairClassification 45.3 39.1 35.5 45.7 50.1
-Reranking 37.5 37.4 35.3 39.5 43.1
-STS 39.9 24.1 18.2 39.9 47.5
-Summarization 28.4 20.9 21.1 29.8 30.8
-Average 34.3 29.9 27.8 35.2 39.2
-Qwen1.5-MoE-A2.7B Classification 43.8 50.3 25.5 47.7 51.2
-Clustering 13.6 27.4 23.2 25.2 26.7
-PairClassification 45.9 46.9 43.4 51.5 53.1
-Reranking 39.6 45.3 41.6 48.5 48.2
-STS 38.8 38.0 35.6 51.8 54.8
-Summarization 28.3 13.4 15.1 31.2 29.7
-Average 35.0 36.9 30.7 42.6 43.9
-DeepSeekMoE-16B Classification 43.4 46.6 44.7 44.4 46.8
-Clustering 13.4 18.1 13.5 17.8 21.9
-PairClassification 45.5 40.9 37.1 46.1 51.9
-Reranking 38.5 38.9 35.1 42.2 45.7
-STS 37.7 26.3 23.3 40.2 45.9
-Summarization 24.9 22.0 18.5 24.4 28.4
-Average 33.9 32.1 28.7 35.9 40.1
-Table 2: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks without prompts and models. The best result for each row is highlighted in bold.
+We provide a detailed evaluation of three state-of-the-art SMoE models: OLMoE-1B-7B (Table  7), Qwen1.5-MoE-A2.7B (Table 8), and DeepSeekMoE-16B (Table  9).
+Our results demonstrate the effectiveness of our method across various models and prompts, comparing its performance against baseline approaches such as Token Choice (TC) and Expert Choice (EC).
+Category Model Dataset Setting Router TC EC MoEE USMoE
+Classification OLMoE Emotion None 24.1 24.5 26.3 25.1 35.8
+Prompt 27.6 49.9 49.0 44.5 54.8
+Toxic None 51.9 58.9 59.9 51.9 62.8
+Prompt 52.3 65.2 61.3 53.4 67.6
+Tweet None 47.7 46.8 48.3 48.4 55.2
+Prompt 49.5 58.0 58.4 57.2 61.7
+Clustering OLMoE Medrxiv None 15.0 17.6 14.8 17.4 20.6
+Prompt 15.8 23.9 27.7 22.0 28.2
+20Groups None 12.4 11.8 9.2 11.5 21.3
+Prompt 16.7 25.7 26.2 24.4 36.0
+Pair Classification OLMoE SemEval None 43.6 35.8 31.3 43.6 50.2
+Prompt 45.7 46.7 40.9 53.8 55.4
+URLCorpus None 47.0 42.4 39.7 47.8 73.2
+Prompt 61.4 77.4 76.9 78.2 82.3
+Reranking OLMoE Ask None 41.3 41.0 39.0 41.4 47.1
+Prompt 43.4 51.9 49.9 50.2 51.6
+SciDocs None 45.5 46.3 46.9 50.8 59.1
+Prompt 53.6 69.6 73.1 75.1 77.2
+StackOver None 25.8 24.8 20.1 26.4 33.2
+Prompt 28.1 32.5 30.0 34.3 36.6
+STS OLMoE Biosses None 39.3 13.6 7.7 29.7 64.0
+Prompt 51.2 61.8 67.6 70.2 75.2
+SickR None 50.3 46.3 26.4 53.0 58.9
+Prompt 51.9 65.7 37.6 66.1 66.5
+STS12 None 40.1 8.6 11.1 37.8 58.2
+Prompt 51.3 53.8 37.5 63.6 66.4
+STS13 None 40.5 21.1 18.2 43.4 61.5
+Prompt 52.5 66.5 40.4 72.7 76.4
+STS14 None 29.5 13.4 13.3 31.7 52.9
+Prompt 41.1 56.8 33.9 64.2 68.2
+STS15 None 30.8 27.8 22.5 33.3 63.2
+Prompt 46.4 69.3 38.4 66.4 72.5
+STS16 None 46.5 38.9 28.9 45.8 60.7
+Prompt 52.4 70.1 49.4 68.3 71.8
+STSBen None 42.2 23.4 17.5 44.5 57.4
+Prompt 48.6 63.6 48.9 70.7 72.1
+Summarization OLMoE Medrxiv None 28.4 20.9 21.1 29.8 31.8
+Prompt 25.6 28.9 29.7 30.4 30.5
+Table 7: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB tasks using the OLMoE model. The best result for each row is highlighted in bold.
+Category Model Dataset Setting Router TC EC MoEE USMoE
+Classification Qwen1.5-MoE-A2.7B Emotion None 27.2 33.9 30.1 34.3 37.1
+Prompt 37.0 48.5 47.4 47.2 51.3
+Toxic None 53.0 61.1 21.3 52.9 61.4
+Prompt 53.4 64.5 19.8 54.1 65.5
+Tweet None 51.1 55.9 25.1 55.9 58.2
+Prompt 56.1 61.1 38.6 60.7 62.3
+Clustering Qwen1.5-MoE-A2.7B Medrxiv None 15.3 23.3 21.3 23.0 25.3
+Prompt 14.2 24.6 19.8 21.8 27.9
+20Groups None 12.0 31.5 25.1 27.4 34.4
+Prompt 14.4 43.8 38.6 38.4 47.0
+Pair Classification Qwen1.5-MoE-A2.7B SemEval None 42.0 38.8 34.7 42.5 47.5
+Prompt 47.0 52.4 46.3 52.4 57.7
+URLCorpus None 49.8 54.9 52.1 60.6 64.8
+Prompt 56.7 68.7 65.8 68.2 75.5
+Reranking Qwen1.5-MoE-A2.7B Ask None 43.1 45.8 43.4 47.3 49.1
+Prompt 43.3 48.3 49.1 49.5 52.8
+SciDocs None 49.6 60.6 55.3 67.0 71.6
+Prompt 50.9 60.1 55.8 68.7 73.0
+StackOver None 26.2 29.5 26.2 31.1 38.5
+Prompt 28.8 31.3 30.2 35.2 44.5
+STS Qwen1.5-MoE-A2.7B Biosses None 33.8 32.5 34.7 49.6 66.6
+Prompt 55.1 55.8 48.5 68.4 74.5
+SickR None 51.0 55.5 40.4 61.0 63.5
+Prompt 50.2 59.7 51.1 64.3 69.1
+STS12 None 40.2 16.9 18.6 46.3 52.3
+Prompt 49.3 25.0 31.8 59.2 62.5
+STS13 None 38.1 42.9 44.2 56.7 67.3
+Prompt 53.3 57.5 54.6 73.4 75.6
+STS14 None 28.1 26.5 25.6 45.4 53.7
+Prompt 40.4 38.8 40.7 60.0 64.8
+STS15 None 34.8 40.5 38.4 46.1 56.1
+Prompt 40.7 52.3 54.2 58.8 66.8
+STS16 None 47.6 51.0 48.1 58.1 64.4
+Prompt 51.6 64.2 65.1 65.7 68.7
+STSBen None 37.0 37.7 34.7 50.9 61.6
+Prompt 45.6 47.8 54.5 64.5 70.0
+Summarization Qwen1.5-MoE-A2.7B Medrxiv None 28.3 13.4 15.1 31.2 40.0
+Prompt 27.0 23.0 21.9 27.3 31.0
+Table 8: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks with Qwen1.5-MoE-A2.7B models. The best result for each row is highlighted in bold.
+Figure 7: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across across MTEB Tasks with PromptEOL [28]. The best result for each row is highlighted in bold.
+Category Model Dataset Setting Router TC EC MoEE USMoE
+Classification DeepSeekMoE-16B Emotion None 26.1 27.4 26.5 27.6 31.3
+Prompt 37.9 48.3 46.1 46.4 52.6
+Toxic None 53.3 60.4 58.1 53.1 61.7
+Prompt 53.1 62.4 62.5 53.6 67.5
+Tweet None 51.0 51.9 49.5 52.6 55.2
+Prompt 54.9 58.4 57.5 58.9 61.0
+Clustering DeepSeekMoE-16B Medrxiv None 15.1 23.0 17.3 22.0 25.7
+Prompt 17.0 25.7 20.9 24.0 27.9
+20Groups None 11.7 13.2 9.7 13.7 16.2
+Prompt 18.6 32.3 19.8 33.0 37.6
+Pair Classification DeepSeekMoE-16B SemEval None 44.6 40.2 32.6 43.5 44.6
+Prompt 48.4 47.2 46.6 51.3 55.7
+URLCorpus None 46.4 41.7 41.6 48.6 62.4
+Prompt 66.5 72.4 61.1 75.4 80.0
+Reranking DeepSeekMoE-16B Ask None 41.7 41.1 40.1 42.3 44.9
+Prompt 43.5 43.8 44.7 46.9 50.6
+SciDocs None 48.2 50.6 44.7 57.1 61.9
+Prompt 58.3 65.6 55.3 72.6 72.9
+StackOver None 25.7 24.9 20.4 27.3 30.2
+Prompt 29.7 27.6 22.6 32.3 33.6
+STS DeepSeekMoE-16B Biosses None 29.5 31.7 27.7 26.8 56.6
+Prompt 47.0 40.1 41.5 57.6 67.2
+SickR None 50.4 47.4 29.4 53.1 60.1
+Prompt 56.0 61.9 38.7 65.8 68.0
+STS12 None 44.0 4.3 13.9 45.0 48.6
+Prompt 57.8 31.0 28.4 64.0 65.9
+STS13 None 36.0 28.4 27.5 41.1 50.7
+Prompt 55.3 56.0 41.2 70.9 76.1
+STS14 None 25.4 12.0 13.0 28.2 41.2
+Prompt 44.9 41.0 31.1 58.6 65.1
+STS15 None 34.8 33.9 25.6 38.7 45.0
+Prompt 49.7 46.5 33.0 58.5 64.1
+STS16 None 44.9 34.4 33.1 46.9 55.2
+Prompt 56.7 58.0 44.2 64.5 67.4
+STSBen None 36.6 18.3 15.8 42.1 51.9
+Prompt 54.9 57.7 39.0 67.8 71.0
+Summarization DeepSeekMoE-16B Medrxiv None 24.9 22.0 18.5 24.4 29.9
+Prompt 29.1 24.4 25.7 29.2 30.7
+Table 9: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks with DeepSeekMoE-16B models. The best result for each row is highlighted in bold.
 (a) Performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) on Classification Task.
@@
 Each benchmark is run 10 times, reporting both the mean and standard deviation to highlight the performance and stability of our method compared to the baselines.
-Transformer-XL(20M) Enwik8 Text8 WikiText-103 lm1b
-[…]
-[…]
-[…]
-SMoE-DR 1.56 1.56 58.37 93.17
-XMoE 1.21 1.28 30.34 58.33
-StableMoE 1.20 1.28 29.97 58.25
-[…]
-Table 3: Performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) across multiple datasets, with BPC on the Enwik8 and Text8 test sets, and perplexity on the WikiText-103 and One Billion Word test sets.
-Lower values are better, with the best results highlighted in bold.
-4.3 Pre-training Result
-Base models training.
-Table 3 summarizes the pre-training results across four datasets (enwik8, text8, WikiText-103, and One Billion Words).
-USMoE consistently outperforms the expert choice baseline and token choice routing methods such as XMoE Chi et al.
-(2022a) and StableMoE Dai et al.
-(2022b) on all four datasets.
-The strength of USMoE lies in its inference efficiency, achieved by leveraging fewer experts.
-Notably, on text8, USMoE surpasses SMoE while utilizing only one and a half experts.
-Additionally, it outperforms SMoE-Dropout (which employs two experts) on One Billion Words, lowering perplexity from 93.17 to 57.55 with the same reduced expert count.
-Furthermore, with only one and a half experts, USMoE reduces FLOPs by 14% compared to SMoE and SMoE-Dropout, which rely on two experts, all while maintaining strong performance.
+Model Dataset TC EC USMoE-Sequence USMoE-Batch
+DeepSeekMoE-16B Emotion 27.4 26.5 27.8 27.4
+Toxic 60.4 58.1 60.1 59.2
+Tweet 51.9 49.5 52.5 51.7
+Table 10: Comparison of USMoE, Token Choice (TC), and Expert Choice (EC) on the classification task. Higher values are better, with the best results highlighted in bold.
+Model Dataset $\alpha$
+0.0 0.3 0.5 0.7 0.9 1.0
+DeepSeekMoE-16B Emotion 27.4 27.1 27.8 27.6 27.7 26.5
+Toxic 60.4 60.0 60.1 56.8 57.3 58.1
+Tweet 51.9 53.2 52.5 53.3 52.9 49.5
+Table 11: Performance comparison of DeepSeekMoE-16B across different classification datasets with varying $\alpha$ values. Higher is better; best results are in bold.
+A.4.3 Training from scratch
 Large models training.
@@
 Transformer-XL(420M) WikiText-103 lm1b
-[…]
+Topkk TC EC USMoE TC EC USMoE
 1 31.70 35.52 25.48 58.65 65.43 56.90
@@
 Lower values are better, with the best results highlighted in bold.
-4.4 Fine-tuning Result
 […]
-[…]
-[…]
-[…]
+USMoE (Topkk=2) 7.7620 81.5 40.1 88.5 87.8
+       (Topkk=1.5) 6.6753 83.8 39.6 88.3 83.0
+TC (Topkk=2) SMoE 7.7620 77.1 35.1 84.4 69.2
 SMoE-DR 78.6 34.4 83.5 66.7
@@
 StableMoE 77.7 34.3 83.9 60.8
-[…]
+EC (Topkk=2) 7.7620 81.5 39.3 88.0 75.6
 Table 13: Accuracy performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) after fine-tuned on various datasets. Higher is better, best results are in bold.
+Training from scratch.
+To assess the effectiveness of our method, we compare USMoE with the Token Choice approaches, including SMoE [27], SMoE-Dropout (abbreviated as "SMoE-DR"), XMoE [10], and StableMoE [13], as well as the Expert Choice approach [58] for pre-training and fine-tuning tasks.
+We follow the approach of [7] and use a base Transformer-XL [14] with four decoder layers.
+We train both base and large-scale versions of Transformer-XL on four datasets (Enwik8, Text8, Wikitext-103, and One Billion Words) for 100k iterations, following the implementation in [7].
+Then we fine-tune the pre-trained weights for text classification tasks, including SST-2 [49], SST-5 [49], IMDB [36], and BANKING77 [4].
+More implementation details and additional results are provided in the Appendix A.
 Fine-tuning.
-We report the results of the fine-tuning experiment on the SST-2, SST-5, IMDB, and BANKING77 datasets in Table  5, using Transformer-XL pre-trained on enwik8.
+We report the results of the fine-tuning experiment on the SST-2, SST-5, IMDB, and BANKING77 datasets in Table  13, using Transformer-XL pre-trained on enwik8.
 Overall, USMoE consistently achieves higher accuracy compared to other baselines across all datasets.
 The results demonstrate that our method is not only effective for pre-training tasks but also performs effectively on existing pre-trained models.
-4.5 Ablation Studies
-We investigate the effectiveness and robustness of USMoE to the different hyper-parameter settings.
-4.5.1 Competitive Learning Strategy Comparison
-Model Dataset TC EC USMoE-Sequence USMoE-Batch
-DeepSeekMoE-16B Emotion 27.4 26.5 27.8 27.4
-Toxic 60.4 58.1 60.1 59.2
-Tweet 51.9 49.5 52.5 51.7
-Table 6: Competitive Learning Strategy comparison of USMoE, Token Choice (TC), and Expert Choice (EC) on the classification task. Higher values are better, with the best results highlighted in bold.
-The Unified Competitive Mechanism is implemented using two approaches: (1) a sequence-based method that compares all tokens within a sequence (referred to as "USMoE-Sequence") and (2) a batch-based method that compares all tokens within a batch or mini-batch (referred to as "USMoE-Batch").
-We evaluate both approaches on the Classification task, with results presented in Table 6.
-The findings indicate that both methods outperform the Expert Choice and Token Choice approaches, demonstrating the effectiveness of our method.
-Notably, the sequence-based approach achieves superior performance in the Classification task, as it ensures that no important tokens are missed within a sequence - an assurance that the batch/mini-batch implementation may not always provide.
-4.5.2 Robustness to the controlling factor
-Model Dataset $\alpha$
-0.0 0.3 0.5 0.7 0.9 1.0
-DeepSeekMoE-16B Emotion 27.4 27.1 27.8 27.6 27.7 26.5
-Toxic 60.4 60.0 60.1 56.8 57.3 58.1
-Tweet 51.9 53.2 52.5 53.3 52.9 49.5
-Table 7: Performance comparison of DeepSeekMoE-16B across different classification datasets with varying $\alpha$ values. Higher is better; best results are in bold.
-The Unified Competitive Score ($\alpha$) enables the model to adjust its scoring mechanism, either favoring a diverse set of experts per sequence or distributing experts more evenly across tokens.
-We evaluate the robustness of the controlling factor $\alpha$ on the classification task using the DeepSeekMoE-16B model, with results presented in Table 7.
-When $\alpha=0.0$, the scoring mechanism aligns with Token Choice, while at $\alpha=1.0$, it follows Expert Choice.
-Overall, USMoE demonstrates strong effectiveness within the range of $\alpha\in(0.3,0.7)$, striking a balance between expert diversity and token importance.
-This range provides an optimal trade-off between enforcing SMoE’s competition policy and enhancing traditional approaches for the task, leading to superior overall performance.
-Notably, all tested $\alpha$ configurations outperform the Expert Choice approach ($\alpha=1.0$).
-Summary and hyper-parameter configuration guidelines.
-Our experiments demonstrate that USMoE not only excels in pre-training and transfer learning but also achieves superior generalization compared to traditional approaches, all without requiring additional training.
-Ablation studies in Section 4.5 indicate that USMoE is highly effective within the range of $\alpha\in(0.3,0.7)$.
-This finding provides a practical guideline for efficiently tuning USMoE through cross-validation.
-5 Conclusion
-In this research, we reformulate Token Choice and Expert Choice Sparse Mixture of Experts from a competitive learning perspective, highlighting their limitations.
-Building on this analysis, we introduce Unified Competitive Learning SMoE (USMoE) - a novel and efficient framework that enhances SMoE through a unified competitive learning approach.
-We theoretically prove that USMoE achieves superior expert selection compared to traditional methods, effectively improving expert learning capacity while mitigating expert collapse.
-As a result, USMoE learns more robust expert representations and overcomes the representation collapse issues commonly observed in conventional SMoE training.
-Experiments across both training-free and training-based settings (including pre-training and fine-tuning) demonstrate that USMoE enables more efficient and effective training and inference compared to state-of-the-art routing strategies.
-Limitations
-Our study focuses on enhancing the efficiency and effectiveness of Large Language Models (LLMs) through SMoE.
-Our approach proves effective in both training and non-training settings.
-While the results are promising, our pre-training experiments were constrained by computational resources, limiting us to medium-scale datasets and a base Transformer-XL model.
-Consequently, further empirical evaluation is required to assess the scalability of USMoE beyond 100B parameters and compare it with other SMoE strategies in modern LLMs.
-Ethics Statement
-Despite encouraging results, training large-scale LLMs remains highly resource-intensive, requiring careful management of computational costs.
-Additionally, our study relies on web-sourced data, which is known to contain gender and racial biases, highlighting the need for further efforts to mitigate these issues.
-Lastly, while our work represents a significant step toward advancing LLMs development, it also emphasizes the importance of robust regularization to prevent potential misuse in harmful applications.
-Appendix A Appendix
-Supplementary Material for “Sparse Mixture of Experts as Unified Competitive Learning”
-This document is organized as follows: Appendix A.1 provides the theoretical proof supporting the discussion in Section 3.2.
-Following this, Appendix A.3 offers a detailed analysis of the SMoE router and explains why our method outperforms the baseline approaches.
-Appendix A.2 presents supplementary experimental results, and Appendix A.4 describes the implementation details in full.
-A.1 Theoretical Proof for Section  3.2
-Let $x\in\mathbb{R}^{N\times d}$ represent an input token embedding and $\mathcal{E}\in\mathbb{R}^{d\times n}$ represent the expert embeddings, where $d$ is the dimension of the SMoE model and $n$ is the number of experts.
-The similarity score $S$ between the input token and the expert embeddings is computed using a dot product:
-$S=x\cdot\mathcal{E}.$ (12)
-Since the softmax function is monotonic, the selection of top-$k$ values remains unchanged under softmax transformation. Formally, we have:
-$\operatorname{TopK}(\operatorname{softmax}(S),k)=\operatorname{TopK}(S,k).$ (13)
-By setting $k=1$, the top similarity score for a given token using the Token Choice approach is:
-$S^{t}_{j}=\max(S_{ik}),\quad\forall k\in[1,n].$ (14)
-Since this operation is performed independently for each input token $i$, we extend this to all tokens $q$ in the batch:
-$S^{t}_{j}\leq\max(S_{qk}),\quad\forall k\in[1,n],\quad\forall q\in[1,N].$ (15)
-By definition, the Unified Competitive Mechanism approach selects the highest similarity score across all tokens and experts:
-$S^{u}_{j}=\max(S_{qk}),\quad\forall k\in[1,n],\quad\forall q\in[1,N].$ (16)
-Thus, we establish the first inequality:
-$S^{t}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N].$ (17)
-Similarly, in the Expert Choice approach, each expert selects the best matching token, leading to:
-$S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N].$ (18)
-Therefore, we conclude:
-$S^{t}_{i}\leq S^{u}_{i},\quad S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,N].$ (19)
-This completes the proof. $\square$
-A.2 Additional Results
-We provide a detailed evaluation of three state-of-the-art SMoE models: OLMoE-1B-7B (Table  8), Qwen1.5-MoE-A2.7B (Table 9), and DeepSeekMoE-16B (Table  10).
-Our results demonstrate the effectiveness of our method across various models and prompts, comparing its performance against baseline approaches such as Token Choice (TC) and Expert Choice (EC).
-Category Model Dataset Setting Router TC EC MoEE USMoE
-Classification OLMoE-1B-7B Emotion None 24.1 24.5 26.3 25.1 28.2
-Prompt 27.6 49.9 49.0 44.5 51.7
-Toxic None 51.9 58.9 59.9 51.9 60.6
-Prompt 52.3 65.2 61.3 53.4 66.6
-Tweet None 47.7 46.8 48.3 48.4 52.1
-Prompt 49.5 58.0 58.4 57.2 59.2
-Clustering OLMoE-1B-7B Medrxiv None 15.0 17.6 14.8 17.4 21.2
-Prompt 15.8 23.9 27.7 22.0 26.3
-20Groups None 12.4 11.8 9.2 11.5 12.6
-Prompt 16.7 25.7 26.2 24.4 34.6
-Pair Classification OLMoE-1B-7B SemEval None 43.6 35.8 31.3 43.6 45.7
-Prompt 45.7 46.7 40.9 53.8 54.2
-URLCorpus None 47.0 42.4 39.7 47.8 54.5
-Prompt 61.4 77.4 76.9 78.2 79.5
-Reranking OLMoE-1B-7B Ask None 41.3 41.0 39.0 41.4 42.4
-Prompt 43.4 51.9 49.9 50.2 51.8
-SciDocs None 45.5 46.3 46.9 50.8 59.9
-Prompt 53.6 69.6 73.1 75.1 76.6
-StackOver None 25.8 24.8 20.1 26.4 27.0
-Prompt 28.1 32.5 30.0 34.3 35.6
-STS OLMoE-1B-7B Biosses None 39.3 13.6 7.7 29.7 45.4
-Prompt 51.2 61.8 67.6 70.2 74.3
-SickR None 50.3 46.3 26.4 53.0 55.1
-Prompt 51.9 65.7 37.6 66.1 67.9
-STS12 None 40.1 8.6 11.1 37.8 44.6
-Prompt 51.3 53.8 37.5 63.6 67.1
-STS13 None 40.5 21.1 18.2 43.4 49.8
-Prompt 52.5 66.5 40.4 72.7 76.2
-STS14 None 29.5 13.4 13.3 31.7 38.7
-Prompt 41.1 56.8 33.9 64.2 67.3
-STS15 None 30.8 27.8 22.5 33.3 43.3
-Prompt 46.4 69.3 38.4 66.4 71.7
-STS16 None 46.5 38.9 28.9 45.8 51.7
-Prompt 52.4 70.1 49.4 68.3 72.7
-STSBen None 42.2 23.4 17.5 44.5 51.1
-Prompt 48.6 63.6 48.9 70.7 71.4
-Summarization OLMoE-1B-7B Medrxiv None 28.4 20.9 21.1 29.8 30.8
-Prompt 25.6 28.9 29.7 30.4 30.9
-Table 8: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks with OLMoE-1B-7B models. The best result for each row is highlighted in bold.
-Category Model Dataset Setting Router TC EC MoEE USMoE
-Classification Qwen1.5-MoE-A2.7B Emotion None 27.2 33.9 30.1 34.3 35.7
-Prompt 37.0 48.5 47.4 47.2 49.9
-Toxic None 53.0 61.1 21.3 52.9 61.4
-Prompt 53.4 64.5 19.8 54.1 66.2
-Tweet None 51.1 55.9 25.1 55.9 56.5
-Prompt 56.1 61.1 38.6 60.7 61.5
-Clustering Qwen1.5-MoE-A2.7B Medrxiv None 15.3 23.3 21.3 23.0 23.4
-Prompt 14.2 24.6 19.8 21.8 25.1
-20Groups None 12.0 31.5 25.1 27.4 30.0
-Prompt 14.4 43.8 38.6 38.4 46.4
-Pair Classification Qwen1.5-MoE-A2.7B SemEval None 42.0 38.8 34.7 42.5 44.2
-Prompt 47.0 52.4 46.3 52.4 54.1
-URLCorpus None 49.8 54.9 52.1 60.6 62.1
-Prompt 56.7 68.7 65.8 68.2 72.7
-Reranking Qwen1.5-MoE-A2.7B Ask None 43.1 45.8 43.4 47.3 47.5
-Prompt 43.3 48.3 49.1 49.5 51.4
-SciDocs None 49.6 60.6 55.3 67.0 65.6
-Prompt 50.9 60.1 55.8 68.7 73.0
-StackOver None 26.2 29.5 26.2 31.1 31.4
-Prompt 28.8 31.3 30.2 35.2 36.6
-STS Qwen1.5-MoE-A2.7B Biosses None 33.8 32.5 34.7 49.6 52.6
-Prompt 55.1 55.8 48.5 68.4 66.2
-SickR None 51.0 55.5 40.4 61.0 63.6
-Prompt 50.2 59.7 51.1 64.3 66.3
-STS12 None 40.2 16.9 18.6 46.3 48.3
-Prompt 49.3 25.0 31.8 59.2 61.4
-STS13 None 38.1 42.9 44.2 56.7 61.8
-Prompt 53.3 57.5 54.6 73.4 75.7
-STS14 None 28.1 26.5 25.6 45.4 48.9
-Prompt 40.4 38.8 40.7 60.0 62.7
-STS15 None 34.8 40.5 38.4 46.1 48.0
-Prompt 40.7 52.3 54.2 58.8 62.5
-STS16 None 47.6 51.0 48.1 58.1 59.6
-Prompt 51.6 64.2 65.1 65.7 68.2
-STSBen None 37.0 37.7 34.7 50.9 55.5
-Prompt 45.6 47.8 54.5 64.5 67.3
-Summarization Qwen1.5-MoE-A2.7B Medrxiv None 28.3 13.4 15.1 31.2 29.7
-Prompt 27.0 23.0 21.9 27.3 48.4
-Table 9: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks with Qwen1.5-MoE-A2.7B models. The best result for each row is highlighted in bold.
-Category Model Dataset Setting Router TC EC MoEE USMoE
-Classification DeepSeekMoE-16B Emotion None 26.1 27.4 26.5 27.6 27.8
-Prompt 37.9 48.3 46.1 46.4 48.8
-Toxic None 53.3 60.4 58.1 53.1 60.1
-Prompt 53.1 62.4 62.5 53.6 64.2
-Tweet None 51.0 51.9 49.5 52.6 52.5
-Prompt 54.9 58.4 57.5 58.9 58.9
-Clustering DeepSeekMoE-16B Medrxiv None 15.1 23.0 17.3 22.0 25.0
-Prompt 17.0 25.7 20.9 24.0 25.8
-20Groups None 11.7 13.2 9.7 13.7 18.7
-Prompt 18.6 32.3 19.8 33.0 37.9
-Pair Classification DeepSeekMoE-16B SemEval None 44.6 40.2 32.6 43.5 46.0
-Prompt 48.4 47.2 46.6 51.3 54.6
-URLCorpus None 46.4 41.7 41.6 48.6 57.9
-Prompt 66.5 72.4 61.1 75.4 75.9
-Reranking DeepSeekMoE-16B Ask None 41.7 41.1 40.1 42.3 45.0
-Prompt 43.5 43.8 44.7 46.9 49.9
-SciDocs None 48.2 50.6 44.7 57.1 61.9
-Prompt 58.3 65.6 55.3 72.6 72.9
-StackOver None 25.7 24.9 20.4 27.3 30.2
-Prompt 29.7 27.6 22.6 32.3 33.6
-STS DeepSeekMoE-16B Biosses None 29.5 31.7 27.7 26.8 35.4
-Prompt 47.0 40.1 41.5 57.6 55.3
-SickR None 50.4 47.4 29.4 53.1 56.7
-Prompt 56.0 61.9 38.7 65.8 67.7
-STS12 None 44.0 4.3 13.9 45.0 46.9
-Prompt 57.8 31.0 28.4 64.0 64.2
-STS13 None 36.0 28.4 27.5 41.1 49.1
-Prompt 55.3 56.0 41.2 70.9 75.2
-STS14 None 25.4 12.0 13.0 28.2 37.3
-Prompt 44.9 41.0 31.1 58.6 63.8
-STS15 None 34.8 33.9 25.6 38.7 40.7
-Prompt 49.7 46.5 33.0 58.5 63.0
-STS16 None 44.9 34.4 33.1 46.9 55.7
-Prompt 56.7 58.0 44.2 64.5 69.7
-STSBen None 36.6 18.3 15.8 42.1 45.6
-Prompt 54.9 57.7 39.0 67.8 69.2
-Summarization DeepSeekMoE-16B Medrxiv None 24.9 22.0 18.5 24.4 28.4
-Prompt 29.1 24.4 25.7 29.2 30.9
-Table 10: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks with DeepSeekMoE-16B models. The best result for each row is highlighted in bold.
-A.3 In-depth Analysis
+A.5 In-depth Analysis
 We visualize the router behavior of USMoE in Figure 11 and contrast it with the router behaviors of the Token Choice approach (Figure 9) and the Expert Choice approach (Figure 10).
@@
 This bias limits the model’s adaptability and effectiveness for downstream tasks.
-USMoE tackles this challenge by introducing the Unified Competitive Mechanism, which promotes more balanced and diverse expert selections, as illustrated in Figure 6.
+USMoE tackles this challenge by introducing the Unified Mechanism, which promotes more balanced and diverse expert selections, as illustrated in Figure 11.
 This enhancement enables USMoE to outperform the baselines on the Emotion Classification task.
@@
 Figure 12: Comparison of the number of experts per sequence for USMoE, Token Choice (TC), and Expert Choice (EC) on the Emotion dataset using the OLMoE-1B-7B model, along with a comparison of token dropping rates for USMoE, TC, and EC during pre-training on the One Billion Word dataset.
-A.4 Implementation Details
-For the Without Training experiments, we implement our method based on the publicly available MoEE implementation Li and Zhou (2024) [^1] https://github.com/tianyi-lab/MoE-Embedding.
+A.6 Implementation Details
+For the Without Training experiments, we implement our method based on the publicly available MoEE implementation [34] [^1] https://github.com/tianyi-lab/MoE-Embedding.
 Due to resource constraints, we validate our method and the baselines using 4-bit quantization with a batch size of 128.
 For the OLMoE-1B-7B model, we conduct experiments on a single H100 GPU, while for the Qwen1.5-MoE-A2.7B and DeepSeekMoE-16B models, we utilize two H100 GPUs.
-The base Transformer-XL variant (Chen et al., 2023b) comprises four Transformer decoder layers, each with an input dimension of 256.
+The base Transformer-XL variant [7] comprises four Transformer decoder layers, each with an input dimension of 256.
 Each layer includes a self-attention mechanism with eight attention heads, followed by a Feed-forward Neural Network (FFN) that has an inner dimension of 512.
@@
 For the larger variants, we scale the model up to twelve layers.
-Our experiments are based on the publicly available SMoE-Dropout implementation (Chen et al., 2023b) [^2] https://github.com/VITA-Group/Random-MoE-as-Dropout.
+Our experiments are based on the publicly available SMoE-Dropout implementation [7] [^2] https://github.com/VITA-Group/Random-MoE-as-Dropout.
 The pre-training experiments were conducted using a single H100 GPU, while the fine-tuning experiments were performed on a single A100 GPU.
 It is important to note that parallel training on multiple GPUs may produce different results.
-A.4.1 Pre-training Experiments
+A.6.1 Pre-training Experiments
+Transformer-XL(20M) Enwik8 Text8 WikiText-103 lm1b
+USMoE (Topkk=2) 1.18 1.20 29.20 56.90
+       (Topkk=1.5) 1.19 1.28 30.67 57.55
+TC (Topkk=2) SMoE 1.20 1.29 30.16 58.00
+SMoE-DR 1.56 1.56 58.37 93.17
+XMoE 1.21 1.28 30.34 58.33
+StableMoE 1.20 1.28 29.97 58.25
+EC (Topkk=2) 1.18 1.24 29.83 58.60
+Table 14: Performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) across multiple datasets, with BPC on the Enwik8 and Text8 test sets, and perplexity on the WikiText-103 and One Billion Word test sets.
+Lower values are better, with the best results highlighted in bold.
 We provide the USMoE implementation details for pre-training our Transformer-XL base and large on enwik8, text8, WikiText-103, and One Billion Word in Table 15.
@@
 One Billion Word 512 11 Adam 2.5e-4 100k
-A.4.2 Fine-tuning Experiments
+A.6.2 Fine-tuning Experiments
 To perform the fine-tuning experiments, we utilize the same model architecture as in the pre-training phase.
@@
 BANKING77 512 16 Adam 1e-4 15
+Limitations
+Our study focuses on enhancing the efficiency and effectiveness of Large Language Models (LLMs) through SMoE.
+Our approach proves effective in both training and non-training settings.
+While the results are promising, our pre-training experiments were constrained by computational resources, limiting us to medium-scale datasets and a base Transformer-XL model.
+Consequently, further empirical evaluation is required to assess the scalability of USMoE beyond 100B parameters and compare it with other SMoE strategies in modern LLMs.
+Ethics Statement
+Despite encouraging results, training large-scale LLMs remains highly resource-intensive, requiring careful management of computational costs.
+Additionally, our study relies on web-sourced data, which is known to contain gender and racial biases, highlighting the need for further efforts to mitigate these issues.
+Lastly, while our work represents a significant step toward advancing LLMs development, it also emphasizes the importance of robust regularization to prevent potential misuse in harmful applications.